### v1

# Geometric Bounds for Persistence

Abstract
In this paper, we bring a new perspective to persistent homology by incorporating key concepts from metric geometry. For a given compact subset $\mathcal{X}$ of a Banach space $\mathbf{Y}$, we study the topological features appearing in family $\mathcal{N}_{\bullet}(\mathcal{X}\subset\mathbf{Y})$ of nested of neighborhoods of $\mathcal{X}$ in $\mathbf{Y}$ and give several geometric bounds on their persistence (lifespans).
First, we discuss the lifespans of these homology classes in terms of their filling radii in $\mathbf{Y}$. By using this, we relate these lifespans to key invariants in metric geometry such as Urysohn width. Next, we bound these lifespans via $\ell^{\infty}$-principal components of the set $\mathcal{X}$, which are also known as Kolmogorov widths. Furthermore, we introduce and study the notion of extinction time of a space $\mathcal{X}$: the critical threshold after which there are no homological features alive in any degree. We also describe novel approaches on how to estimate Čech and Vietoris–Rips extinction times of a set $\mathcal{X}$ by relating $\mathcal{X}$ to its über-contractible cores and to its tight span, respectively.
1 Introduction
In the past decade, several approaches within Topological Data Analysis (TDA) have been developed in order to facilitate the discovery of patterns in various forms of data. Persistent Homology (PH) – one of the the key concepts in TDA – provides an effective multiscale feature extraction method. In particular, PH has become quite popular in various machine learning tasks, ranging from bioinformatics to finance, materials science to network analysis  [${}^{+}$18, HMR21, ${}^{+}$20, SL22, Giu22, PLX22].
In this paper, we study some aspects of the PH methodology with tools from metric geometry. Our results provide novel quantitative interpretations of the PH output.
[…]
In many applications, the lifespan of a topological feature $\omega$ has crucial significance as it is considered to reflect the size or degree of “importance” of $\omega$. Indeed, in practice, the topological features with long lifespans are thought to represent the main shape characteristics of a dataset whereas the ones with short lifespans are usually regarded as (topological) noise. Therefore, determining and interpreting the lifespans of these topological features appearing in the persistence diagram has critical significance for applications.
In this paper, we aim to give a geometric interpretation of these lifespans by relating them to several notions from metric geometry. We study the lifespans of individual homology classes $\omega$ appearing throughout the Vietoris-Rips and Čech filtrations by resorting to the notion of filling radius and to several notions of width. Our results show that the lifespans of these homology classes are determined by the geometric size of their representatives in the filtration. We first discuss implications of the the absolute (Gromov’s) and relative filling radius of a homology class (2.30) in our setting. Then, we observe that the Čech lifespan of a homology class $\omega$ is equal to its relative filling radius in its ambient space (Section 2.3).  Next, we give several bounds for the lifespans of individual homology classes by resorting to the notions of Urysohn width, Alexandrov width, and Kolmogorov width. Since all these notions of width are monotonically non-increasing with respect to the dimension parameter (e.g., $\operatorname{UW}_{k}(\mathcal{X})\geq\operatorname{UW}_{k+1}(\mathcal{X})$), any estimate for degree-$k$ automatically applies to homology classes in higher degrees.
{restatable*}
[VR Lifespans via Urysohn Width]corcorvruw Let $\mathcal{X}$ be a compact metric space, and let $\omega\in\mathrm{Spec}_{k}(\mathcal{V}_{\bullet}(\mathcal{X}))$, $k\geq 1$. Then,
$d_{\omega}-b_{\omega}\leq\operatorname{UW}_{k-1}(\overline{\mathcal{N}}_{b_{\omega}}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))).$
In particular, $d_{\omega}-b_{\omega}\leq\operatorname{UW}_{k-1}(\operatorname{E}(\mathcal{X}))$.
Here, $\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})$ denotes the open $r$-neighborhood of $\mathcal{X}$ in $\mathbf{Y}$ and $\operatorname{E}(\mathcal{X})$ denotes the tight span of $\mathcal{X}$. Notice that in a special case, if $\mathcal{X}$ is a closed $k$-manifold, and $\omega=[\mathcal{X}]$ is the the fundamental class corresponding of $\mathcal{X}$, since $b_{\omega}=0$ in that case, the result above implies that Vietoris-Rips lifespan of $\omega$ is bounded above by the Urysohn width of $\mathcal{X}$, i.e., $d_{\omega}\leq\operatorname{UW}_{k-1}(\mathcal{X})$ (this particular bound goes back to Gromov; it follows from [Gro83, Appendix 1, Example after Lemma (B)] combined with [Gro83, Appendix 1, Proposition (D)]).
Next, we give several bounds for Čech lifespans. The first one is via Alexandrov widths.
{restatable*}
[Čech Lifespans via Alexandrov Width]corcoraw Let $\mathcal{X}$ be compact subset of a Banach space $\mathbf{Y}$ and let $\omega\in\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$, $k\geq 1$. Then,
$d_{\omega}-b_{\omega}\leq\operatorname{AW}_{k-1}(\overline{\mathcal{N}}_{b_{\omega}}(\mathcal{X}\subset\overline{\mathrm{conv}}(\mathcal{X}))\subset\mathbf{Y}).$
In particular, $d_{\omega}-b_{\omega}\leq\operatorname{AW}_{k-1}(\overline{\mathrm{conv}}(\mathcal{X})\subset\mathbf{Y})$.
Here $\overline{\mathrm{conv}}(\mathcal{X})\subset\mathbf{Y}$ denotes the closure of the convex hull of $\mathcal{X}\subset\mathbf{Y}$.
Further, we introduce a new notion of width, called treewidth (3.9), in order to obtain a finer estimate on Čech lifespans (3.12). Notice that in the estimates above, for each class $\omega$, in order to get an upper bound for its lifespan, the corresponding width needs to be calculated for $\mathcal{N}_{b_{\omega}}(\mathcal{X}\subset\cdot)$, not $\mathcal{X}$ itself. In the following, we get rid of this dependency. The notion of treewidth permits establishing a certain multiplicative bound on lifespans.
{restatable*}
[Čech Lifespans via Treewidth]corcortw Let $\mathcal{X}$ be compact subset in a Banach space $\mathbf{Y}$ and let $\omega\in\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$, with the birth time $b_{\omega}\geq 1$. Then,
$\dfrac{d_{\omega}}{b_{\omega}}\leq C+1+\operatorname{TW}^{C}_{k}(\mathcal{X}\subset\mathbf{Y}).$
Next, we consider the $\ell^{\infty}$-version of principal component analysis (${}_{\infty}$) for a compact subset $\mathcal{X}$ of a Banach space $\mathbf{Y}$ (e.g., a point cloud in $\mathbb{R}^{N}$). By using the estimates in previous sections, we relate the $(k+1)$-variance $\nu_{k+1}(\mathcal{X}\subset\mathbf{Y})$ with the lifespans of classes $\omega$ appearing throughout the Čech filtration (Section 3.5). We note that the variance $\nu_{k+1}(\mathcal{X}\subset\mathbf{Y})$ is also known as the $k^{th}$ Kolmogorov width of $\mathcal{X}$ in the approximation theory literature (3.24).
{restatable*}
[Čech lifespans via $\infty$-Variance]corcorpca Let $\mathcal{X}$ be a compact subset of a Banach space $\mathbf{Y}$ and let $\omega\in\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$, $k\geq 0$. Then,
$d_{\omega}-b_{\omega}\leq\nu_{k+1}(\mathcal{X}\subset\mathbf{Y})=\operatorname{KW}_{k}(\mathcal{X}\subset\mathbf{Y}).$
Note that by monotonicity of widths (e.g., $\operatorname{AW}_{k}(\mathcal{X}\subset\mathbf{Y})\geq\operatorname{AW}_{k+1}(\mathcal{X}\subset\mathbf{Y}))$, all these bounds apply to homology classes of degree higher than $k$, too. While these lifespan bounds depend on the homology degree, we next give a general bound for Čech lifespans which is independent of degree. We achieve this by generalizing Katz’s notion of spread to the extrinsic setting and introducing the notion that we call überspread (Section 4). Überspread basically measures the Hausdorff distance of the space to the closest übercontractible space (a contractible space where all neighborhoods are also contractible) in an ambient space. With this notion, we generalize the existing VR-lifespan estimates with spread in [LMO23] to Čech-lifespans in any degree.
{restatable*}
[Čech Lifespans via Überspread]thmthmuberspread Let $\mathcal{X}$ be a compact subset of a Banach space $(\mathbf{Y},\|\cdot\|)$. Let $\omega\in\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$ for any $k\geq 0$. Then,
$d_{\omega}-b_{\omega}\leq 2\operatorname{{\ddot{u}}-spread}(\mathcal{X}\subset\mathbf{Y}).$
While the results above are very effective for bounding individual lifespans ($d_{\omega}-b_{\omega}$), we next attack a more general question: how to obtain a global bound for death times of homology classes across all degrees? To do this we introduce a notion, called extinction time, representing the maximal threshold after which there is no nontrivial homology class in any degree $k\geq 0$. We bound both Čech extinction times $\widecheck{\xi}(\mathcal{X}\subset\mathbf{Y})$ and Vietoris–Rips extinction times $\xi(\mathcal{X})$ by relating $\mathcal{X}$ to their convex hulls and tight spans, respectively. In the case of Čech filtrations, we introduce a notion called convexity deficiency, $\operatorname{cdef}(\mathcal{X}\subset\mathbf{Y})$, which is the Hausdorff distance of a space $\mathcal{X}$ to its convex hull in $\mathbf{Y}$ (Section 5.1).
{restatable*}
[Bounding Čech Extinction]thmthmcechextintion Let $\mathcal{X}$ be a compact subset of a Banach space $(\mathbf{Y},\|\cdot\|)$. Then,
$\widecheck{\xi}(\mathcal{X}\subset\mathbf{Y})\leq\operatorname{cdef}(\mathcal{X}\subset\mathbf{Y}).$
In the VR case, we define an analogous notion called hyperconvexity deficiency, $\operatorname{hcdef}(\mathcal{X})$, which is the Hausdorff distance between the isometric copy of $\mathcal{X}$ to its tight span $\operatorname{E}(\mathcal{X})$ (Section 5.2). We then show that a result analogous to the above theorem is also true in the VR-case (5.9).
From a conceptual perspective, our results can be considered as relating a given space $\mathcal{X}$ with another space $\Lambda_{\mathcal{X}}$ (its core). Estimates on the lifespans of homological features created in the course of increasing the radius of neighborhoods of $\mathcal{X}$ are obtained through the distance between $\mathcal{X}$ and $\Lambda_{\mathcal{X}}$. In our study of different notions of widths (Section 3), the role of core $\Lambda_{\mathcal{X}}$ is, roughly speaking, played by the $k$-dimensional space which is closest to $\mathcal{X}$ and $k$-width can be regarded as the distance between $\mathcal{X}$ and $\Lambda_{\mathcal{X}}$.In the section where we consider the notion of spread (Section 4), the core $\Lambda_{\mathcal{X}}$ becomes an übercontractible space and the distance between $\mathcal{X}$ and $\Lambda_{\mathcal{X}}$ upper bounds the lifespans in any degree. Here, the condition on the core is having trivial homology groups rather than a condition on its dimension. Similarly, in the section where we consider extinction times (Section 5), we enforce strong geometric conditions (convexity or hyperconvexity) on $\Lambda_{\mathcal{X}}$ again to obtain our extinction bounds through the distance between $\Lambda_{\mathcal{X}}$ and $\mathcal{X}$.
We would like to point out that from one side, while we give upper bounds for quantities from applied topology (e.g., lifespans of homology classes) with metric geometry concepts, at the same time, these results provide computational lower bounds for metric geometry notions based on ideas from persistent homology. This bidirectional relationship underscores the synergy between these fields, enabling us to derive insights that contribute to both domains.
1.1 Related Work
In this work, we aim to build a bridge between two seemingly far fields, i.e., applied algebraic topology and metric geometry. In a way, both applied algebraic topology and metric geometry are interested in similar question with different tools: How to measure the size of a set/space/manifold? In applied algebraic topology, persistent homology is a very effective tool for accomplishing this aim, and the lifespans (or persistence) of topological features induced by Vietoris–Rips (or Čech) filtrations represent the size of the corresponding topological features. In particular, Vietoris–Rips (or Čech) complexes were invented in order to transform a given metric space into a simplicial complex by maintaining its topological information, thus enabling an effective cohomology theory for metric spaces [Vie27, Hau95, Bor48]. There are several works in the literature which study Vietoris–Rips filtrations in different settings [${}^{+}$09, AA17, AAR19, ${}^{+}$20, AMMW22, CDSO14, LMO23, Vir20, Vir22].
On the other hand, from the metric geometry side, estimating the size of a manifold has been a key problem for years. Gromov introduced and studied the notion of filling radius in his seminal paper [Gro83], and several other notions of “largeness” in [Gro86]. Before Gromov, certain relative, or extrinsic, versions of the filling radius were studied in geometric analysis in connection with the isoperimetric inequality [BS83, FF60, MS73]. Gromov brought the filling radius to the realm of systolic geometry, and the study of scalar curvature [GL83]. Several other measures of size of a given manifold or metric space, now known as widths, were also studied and popularized by Gromov [Gro83, Gro88]. The interplay between filling radii, widths, and other metric invariants (from volume and beyond) has been an active research area since then. Katz determined the filling radius of spheres and other essential spaces [Kat83, Kat89, Kat90, Kat91]. Several authors have studied the filling radii in comparison with other largeness properties [Cai94, BH10]. Guth proved some related conjectures of Gromov [Gut11, Gut17]. Sabourau, Nabutovsky, and Rotman related the filling radius with sweepouts of manifolds [Sab20, NRS21]. Bounds on filling radius in terms of Hausdorff content follow from very general isoperimetric estimates due to Liokumovich, Lishak, Nabutovsky, and Rotman [LLNR22].
In the past couple of years, there have been some articles exploring connections between these two worlds working on analogous problems with different tools. With this aim, Lim, Mémoli, and Okutan related the filling radius of a closed manifold to the bar corresponding to fundamental class in the top VR persistence diagram [LMO23]. As shown in [LMO23, Section 9.3.2], the stability of persistence diagrams of Vietoris-Rips filtrations can be used to obtain lower bounds for the Gromov-Hausdorff distance between spheres which arise from considerations related to their filling radii, which in [LMS22] were shown to not be tight. The polymath project  [${}^{+}$22] recently furthered this line work. Recently, Adams–Coskunuzer used a well-known quantity in metric geometry, Urysohn width, to estimate the lifespans in the persistence diagram of a given space [AC22]. In [Vir22], Virk studied the relation between higher persistence diagrams for large degrees and lower dimensional features in the manifold setting.
In this paper, we aim to establish a direct connection between notions from applied algebraic topology and notions in metric geometry by relating several quantities used to measure the size of a manifold in both subfields.
Ackowledgements
This work was supported by the following grants from the National Science Foundation (Grants # DMS-1926686, DMS-2202584, DMS-2220613, DMS-2229417, DMS-1723003, IIS-1901360, CCF-1740761, and DMS-1547357), from the Simons Foundation (Grant # 579977) and from the BSF (under grant # 2020124). The project has received funding from the HORIZON EUROPE Research and Innovation programme under Marie Skłodowska-Curie grant agreement number 101107896. A major part of this work was done while the first-named author was a postdoc at the Institute for Advanced Study, Princeton.
2 Background
In this section, we overview the basic notions from applied algebraic topology and metric geometry used in the remainder of the paper. We give a summary of our notations in Table 1 in the Appendix.
2.1 Persistent Homology
Persistent Homology (PH) is a methodology with roots in Applied Algebraic Topology which helps capturing different characteristics of a given metric space. In the past two decades, PH has been employed as a powerful mathematical machinery for discovering patterns in data in applications within Machine Learning and Data Science. For more details on PH and its use in other settings, see [Car09, Ede13, CM21, DW22].
Neighborhood Notation: Throughout the paper, we use both open and closed neighborhoods and adopt the following notation. When $\mathcal{X}$ is a subset of a metric space $\mathcal{Z}$, through $\mathcal{N}_{r}(\mathcal{X}\subset\mathcal{Z})$ we will denote the open $r$-neighborhood of $\mathcal{X}$ in $\mathcal{Z}$ while $\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\mathcal{Z})$ will denote the closed $r$-neighborhood of $\mathcal{X}$ in $\mathcal{Z}$.
2.1.1 Filtrations
As mentioned in the previous section, one needs a filtration (a nested family of topological spaces or abstract simplicial complexes) $\Delta_{\bullet}$ in order to employ the PH machinery. One of the most natural examples of such filtrations arises by considering $r$-neighborhoods, i.e., for $\mathcal{X}$ a subspace of a metric space $\mathcal{Z}$, then the family $\{\mathcal{N}_{r}(\mathcal{X}\subset\mathcal{Z})\}_{r\geq 0}$ defines a filtration.  While there are several ways to induce such families, from a computational perspective, a tractable method is to induce a family of simplicial complexes from a given dataset. In this paper, we are only interested in the most common ones, Vietoris–Rips complexes and Čech complexes. Note that most of our results for Čech complexes directly generalize to the alpha-complexes naturally (2.9).
Definition 2.1 (Vietoris–Rips Complexes).
Let $(\mathcal{X},\mathbf{d}_{\mathcal{X}})$ be a compact metric space. For $r>0$, its Vietoris–Rips complex is the abstract simplicial complex $\mathcal{V}_{r}(\mathcal{X})$ where a $k$-simplex $\sigma=[x_{i_{0}},x_{i_{1}},\dots,x_{i_{k}}]\in\mathcal{V}_{r}(\mathcal{X})$ if and only if $\mathbf{d}_{\mathcal{X}}(x_{i_{m}},x_{i_{n}})<r$ for any $0\leq m,n\leq k$.
In the sequel, given a metric space $Z$, a point $z\in Z$ and $r>0$, by $B_{r}(z)$ we will denote the open ball of radius $r$ around $z$.
Definition 2.2 (Čech Complexes).
Let $\mathcal{X}$ be a compact subset of a Banach space $\mathbf{Y}$. For $r>0$, Čech complex is the abstract simplicial complex $\mathcal{C}_{r}(\mathcal{X}\subset\mathbf{Y})$ where a $k$-simplex $\sigma=[x_{i_{0}},x_{i_{1}},\dots,x_{i_{k}}]\in\mathcal{C}_{r}(\mathcal{X}\subset\mathbf{Y})$ if and only if $\bigcap_{m=0}^{k}B_{r}(x_{i_{m}})\neq\emptyset$ in $\mathbf{Y}$.
Through the geometric realization functor, the nested families of simplicial complexes provided by the Vietoris-Rips and Čech complexes induce filtrations.
Definition 2.3 (Filtration).
A filtration of a topological space $Z$ is a collection $\Delta_{\bullet}=\big{(}\Delta_{r},\iota_{r,s}\big{)}_{0<r\leq s}$ such that for each $0<r\leq s$, $\Delta_{r}$ is a subset of $Z$ and $\iota_{r,s}:\Delta_{r}\hookrightarrow\Delta_{s}$ is the inclusion map.
When there is no risk of confusion we will simply say that $\Delta_{\bullet}$ is a filtration without mentioning the topological space $Z$, with the understanding that $Z$ can be recovered as the colimit of $\Delta_{\bullet}$.
Example 2.4.
One example of filtrations arising in geometric scenarios is the following. Let $\mathcal{X}\subset Z$ be a nonempty compact subset of a metric space $(Z,\mathbf{d}_{Z})$. Then, one considers the filtration $\mathcal{N}_{\bullet}(\mathcal{X}\subset Z)$ given for each $r>0$ by
$\mathcal{N}_{r}(\mathcal{X}\subset Z):=\bigcup_{x\in\mathcal{X}}B_{r}(x),$
the open $r$-neighborhood of $X$ in $Z$. Of particular relevance to this paper will be the case when $Z=\mathbf{Y}$, a Banach space.
In the remainder of the paper, we use the notation $\mathcal{V}_{\bullet}(\mathcal{X})$ and $\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y})$ to denote the filtrations induced by the (geometric realizations of the) Vietoris-Rips and Čech complexes of $\mathcal{X}$, respectively. Similarly, we will denote by $\mathcal{N}_{\bullet}(\mathcal{X}\subset\mathbf{Y})$ the filtration induced by considering subsequent neighborhoods of the compact subset $\mathcal{X}$ inside of the Banach space $\mathbf{Y}$ and will refer to any such filtration arising in that manner as a neighborhood filtration.
Given that the definitions of the Vietoris-Rips and Čech simplicial complexes are similar, one can expect some relationships between these complexes. By direct computation ([EH10]), it is straightforward to see that for any $\mathcal{X}\subset\mathbf{Y}$,
$\mathcal{C}_{r}(\mathcal{X}\subset\mathbf{Y})\subset\mathcal{V}_{2r}(\mathcal{X})\subset\mathcal{C}_{2r}(\mathcal{X}\subset\mathbf{Y}).$
The Nerve Theorem directly relates the $r$-neighborhoods of $\mathcal{X}$ in $\mathbf{Y}$, $\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})$, with the induced Čech simplicial complexes.
Lemma 2.5 (Nerve Theorem [Ale28, Bor48]).
Let $\mathcal{X}$ be a compact subset of a Banach space $\mathbf{Y}$. For any $r>0$, $\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})$ and $\mathcal{C}_{r}(\mathcal{X}\subset\mathbf{Y})$ are homotopy equivalent to each other, i.e.,
$\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})\simeq\mathcal{C}_{r}(\mathcal{X}\subset\mathbf{Y}).$
There is a “persistent”, or “functorial”, version of this result.
Theorem 2.6 (Persistent Nerve Theorem [LMO23, Proposition 4.1]).
There exist homotopy equivalences $\varphi_{s}:\mathcal{C}_{s}(\mathcal{X}\subset\mathbf{Y})\rightarrow\mathcal{N}_{s}(\mathcal{X}\subset\mathbf{Y})$ for each $s>0$ such that for each $t>s>0$ the following diagram commutes up to homotopy:
[…]
Theorem 2.6 implies that the persistent homology of $\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y})$ is isomorphic to that of $\mathcal{N}_{\bullet}(\mathcal{X}\subset\mathbf{Y})$, a fact that we will repeatedly use in the sequel. Since any compact metric space $\mathcal{X}$ can be isometrically embedded into $L^{\infty}(\mathcal{X})$ (endowed with the sup norm between functions) via its Kuratowski embedding (2.23), one obtains an analogous result providing a connection between the Vietoris–Rips filtration and the nested neighborhoods $\{\mathcal{N}_{s}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))\subseteq\mathcal{N}_{t}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))\}_{0<s<t}$ of $\mathcal{X}\subset L^{\infty}(\mathcal{X})$.
Corollary 2.7 ([LMO23, Theorem 5]).
There exist homotopy equivalences $\varphi_{s}:\mathcal{V}_{2s}(\mathcal{X})\rightarrow\mathcal{N}_{s}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))$ for each $s>0$ such that for each $t>s>0$ the following diagram commutes up to homotopy:
[…]
Remark 2.8.
The result above remains valid if $L^{\infty}(\mathcal{X})$ is substituted by any other injective metric space admitting an isometric embedding of $\mathcal{X}$; see Section 2.2 for the definition and Section 5.2, where we in particular utilize the tight span $\operatorname{E}(\mathcal{X})$ as one such choice.
Remark 2.9 (Alpha complexes).
Note that while we only discuss VR and Čech filtered complexes in our paper, our results on Čech lifespans naturally apply to lifespans of homology classes induced by alpha complexes as $\mathcal{A}_{r}(\mathcal{X}\subset\mathbf{Y})\simeq\mathcal{C}_{r}(\mathcal{X}\subset\mathbf{Y})$ where $\mathcal{A}_{r}(\mathcal{X}\subset\mathbf{Y})$ represents the alpha complex induced by $\mathcal{X}$ with distance threshold $r\geq 0$; see [DW22].
2.1.2 Persistent Homology
Here, we recall basic notions pertaining to persistent homology that are necessary for our setting. For further details, see [LMO23, Section 2.1].
Definition 2.10 (Persistence Module).
A persistence module $(V_{r},\Phi_{r,s})_{0<r\leq s}$ over $\mathbb{R}_{>0}$ is a family of $\mathbb{F}$-vector spaces $V_{r}$ for some field $\mathbb{F}$ with morphisms $\Phi_{r,s}:V_{r}\to V_{s}$ for each $r\leq s$ such that
  • $\Phi_{r,r}=\mathrm{id}_{V_{r}}$,
  • $\Phi_{s,t}\circ\Phi_{r,s}=\Phi_{r,t}$ for each $r\leq s\leq t$.
For conciseness we will denote by $V_{\bullet}$ the persistence module given by $(V_{r},\Phi_{r,s})_{0<r\leq s}$. The morphisms $\Phi_{\bullet,\bullet}$ are referred to as the structure maps of $V_{\bullet}$. Note that a persistence module $V_{\bullet}$ can be regarded as a functor from the poset $(\mathbb{R}_{>0},\leq)$ to the category of vector spaces.
Definition 2.11 (Interval persistence module).
Given an interval $I$ in $\mathbb{R}_{>0}$ and a field $\mathbb{F}$, the interval persistence module induced by $I$ is the persistence module $\mathbb{F}_{\bullet}[I]$ is defined as follows: The vector space at $r$ is $\mathbb{F}$ if $r$ is in $I$ and zero otherwise. Given $r\leq s$, the morphism corresponding to the pair $(r,s)$ is the identity if $r,s$ are both contained in $I$ and zero otherwise.
Definition 2.12 (Barcode and Persistence Diagram).
For a given persistence module $V_{\bullet}$, if there is a multiset of intervals $(I_{\lambda})_{\lambda\in\Lambda}$ such that $V_{\bullet}$ is isomorphic to $\bigoplus_{\lambda\in\Lambda}\mathbb{F}_{\bullet}[I_{\lambda}]$, then that multiset is referred to as a (persistence) barcode associated to the persistence module $V_{\bullet}$. Persistence Modules admitting such a multiset of intervals are said to be interval decomposable. The persistence diagram of $V_{\bullet}$ is then given as the multiset of points $(b_{\lambda},d_{\lambda})\in\mathbb{R}^{2}$, where $b_{\lambda}$, is the left endpoint of $I_{\lambda}$ and $d_{\lambda}$ is its right endpoint. [^3] Not every persistence module is interval decomposable; see [CB15].
In applied algebraic topology, many persistence modules arise as follows.
Definition 2.13 (Persistent Homology of a filtration).
For any $k\geq 0$, applying the $k$-dimensional reduced homology functor (with coefficients in a field $\mathbb{F}$) to a filtration $\Delta_{\bullet}=\big{(}\Delta_{r},\iota_{r,s}\big{)}_{0<r\leq s}$ produces the persistence (homology) module $\operatorname{\tilde{H}}_{k}(\Delta_{\bullet},\mathbb{F})=(\operatorname{\tilde{H}}_{k}(\Delta_{r};\mathbb{F}),\Phi^{k}_{r,s})_{0<r\leq s}$ where the morphisms $\Phi^{k}_{r,s}$ are those induced by $\iota_{r,s}$. [^4] Note that we are using reduced homology in our definition in order to dispense with the usual infinite length bar at the level of degree zero persistent homology.
In what follows we will often drop the field $\mathbb{F}$ from the notation. Under suitable assumptions, the persistence modules obtained from filtrations, as described above, are interval decomposable. In particular, the persistence modules obtained from neighborhood filtrations of compact subsets of a Banach space are interval decomposable (so that they admit barcodes).
Theorem 2.14 ([LMO23, Theorem 1]).
Assume $\mathcal{X}$ is a compact subset of a Banach space $\mathbf{Y}$. Then there is a (unique) persistence barcode associated to the persistence module $\operatorname{\tilde{H}}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$. In particular, $\operatorname{\tilde{H}}_{k}(\mathcal{V}_{\bullet}(\mathcal{X}))$ admits (unique) persistence barcode. [^5] In [LMO23, Theorem 1] the authors only contemplate the case of the $\mathbf{Y}$ being equal to $L^{\infty}(\mathcal{X})$ for some compact metric space $\mathcal{X}$. However, the proof of their result directly applies to the setting in the statement.
We will henceforth use $\mathrm{PD}_{k}(\mathcal{X})$ and $\widecheck{\mathrm{PD}}_{k}(\mathcal{X}\subset\mathbf{Y})$ to respectively denote the persistence diagrams of the Vietoris-Rips and Čech filtrations of $\mathcal{X}$.
2.1.3 Stability Theorems
Persistence diagrams are an effective methodology for encoding topological properties of the space $\mathcal{X}$ and its neighborhoods in $\mathbf{Y}$. Persistence diagrams are stable, as expressed by the following stability theorems. Informally, the stability theorem states that if the shape and the size of two spaces are similar, then their persistence diagrams are close to each other. For a formal statement, let $\mathbf{d}_{b}(.,.)$ denote the bottleneck distance between persistence diagrams [EH10]. Let $\mathbf{d}_{\mathrm{H}}^{\mathbf{Y}}$ be the Hausdorff distance between two subsets of the same Banach space $\mathbf{Y}$, and let $\mathbf{d}_{\mathrm{GH}}$ be the Gromov–Hausdorff distance between two metric spaces [BBI01].
Lemma 2.15 (Stability Theorem—VR; [${}^{+}$09, Theorem 3.1] and [CDSO14, Theorem 5.2]).
Let $\mathcal{X}$ and $\mathcal{X}^{\prime}$ be two compact metric spaces. Then,
$\mathbf{d}_{b}(\mathrm{PD}_{k}(\mathcal{X}),\mathrm{PD}_{k}(\mathcal{X}^{\prime}))\leq 2\,\mathbf{d}_{\mathrm{GH}}(\mathcal{X},\mathcal{X}^{\prime}).$
Lemma 2.16 (Stability Theorem—Čech;  [CDSO14, Theorem 5.6]).
Let $\mathcal{X},\mathcal{X}^{\prime}$ be two compact subsets of a Banach space $\mathbf{Y}$. Then,
$\mathbf{d}_{b}(\widecheck{\mathrm{PD}}_{k}(\mathcal{X}\subset\mathbf{Y}),\widecheck{\mathrm{PD}}_{k}(\mathcal{X}^{\prime}\subset\mathbf{Y}))\leq\mathbf{d}^{\mathbf{Y}}_{\mathrm{H}}(\mathcal{X},\mathcal{X}^{\prime}).$
Note that Lemma 2.16 implies Lemma 2.15. Indeed, this was implicity used in the proof of [${}^{+}$09, Theorem 3.1]. Notice that the coefficient $2$ does not appear in the second stability theorem.
2.1.4 Homological Spectra
As described in the introduction, the chief goal of our paper is to provide effective bounds for the lifetime of homology classes that appear along a (geometric) filtration of a metric space. We will formulate and realize this goal in a setting that encompasses, but is more general than, persistence diagrams.
We recall some additional definitions and results from [LMO23].
Definition 2.17.
For an integer $k\geq 0$, a given field $\mathbb{F}$, and a filtration $\Delta_{\bullet}=\big{(}\Delta_{r},\iota_{r,s}\big{)}_{0<r\leq s}$, let
$\mathrm{Spec}_{k}(\Delta_{\bullet}):=\bigcup_{r>0}\bigg{(}\operatorname{\tilde{H}}_{k}(\Delta_{r};\mathbb{F})\backslash\{0\}\times\{r\}\bigg{)}$ (1)
be the $k$-th homological spectrum of $\Delta_{\bullet}$ (with coefficients in $\mathbb{F})$.
Now, fix an arbitrary $(\omega,s)\in\mathrm{Spec}_{k}(\Delta_{\bullet})$. Then, let
$b_{(\omega,s)}:=\inf\{r>0:r\leq s\text{ and }\exists\text{ nonzero }\omega_{r}\in\operatorname{\tilde{H}}_{k}(\Delta_{r})\text{ such that}(\iota_{r,s})_{\ast}(\omega_{r})=\omega\},$ (2)
$d_{(\omega,s)}:=\sup\{t>0:t\geq s\text{ and }\exists\text{ nonzero }\omega_{t}\in\operatorname{\tilde{H}}_{k}(\Delta_{t})\text{ such that}(\iota_{s,t})_{\ast}(\omega)=\omega_{t}\}.$ (3)
Whenever $\Delta_{\bullet}$ is a neighborhood filtration $\mathcal{N}_{\bullet}(\mathcal{X}\subset\mathbf{Y})$, as in [LMO23, Theorem 8], one has that $b_{(\omega,s)}<s\leq d_{(\omega,s)}$. [^6] In general, the type of intervals (open-open,closed-open, etc) one obtains depends on whether the filtration is defined via open or closed neighborhoods. Note that we’ve defined neighborhood filtrations via open neighborhoods. Let
$I_{(\omega,s)}:=(b_{(\omega,s)},d_{(\omega,s)}].$
Informally, the interval $I_{(\omega,s)}$ encodes the maximal region around $s\in\mathbb{R}_{>0}$ inside which the class $\omega$ is “alive”.777 Note that, since we are using reduced homology, there is no degree 0 class $\omega$ such that $d_{(\omega,s)}=\infty$.
Definition 2.18.
The value $b_{(\omega,s)}$ is referred to as the birth time of $\omega$ whereas $d_{(\omega,s)}$ is the death time of $\omega$. The value $d_{(\omega,s)}-b_{(\omega,s)}$ will be referred to as the lifespan of $\omega.$
Remark 2.19.
We focus on the special case of neighborhood filtrations. The birth time $b_{(\omega,s)}$ was defined as the infimum of all times $r\leq s$ when a “predecessor” of $\omega$ exists in $\operatorname{\tilde{H}}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y}))$, and it is natural to ask if there is a homology class that is supported on
$\bigcap\limits_{b_{(\omega,s)}<r\leq s}\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})=\overline{\mathcal{N}}_{b_{(\omega,s)}}(\mathcal{X}\subset\mathbf{Y})$
and also homologous to $\omega$ in $\mathcal{N}_{s}(\mathcal{X}\subset\mathbf{Y})$. It turns out this is not the case even for the neighborhood filtration of a compact set in $\mathbb{R}^{2}$. Namely, there exists a compact set $\mathcal{X}\subset\mathbb{R}^{2}$, known as the Warsaw circle or closed topologist’s sine curve [Bor75], satisfying the following counter-intuitive property: its first singular homology is zero, but every open neighborhood of it contains a homologically non-trivial circle. Every two of those circles are homologous to each other (within the union of the two neighborhoods). But these circles do not converge, as we shrink the neighborhood, to a non-trivial homology class of $\mathcal{X}$, because $\operatorname{\tilde{H}}_{1}(\mathcal{X})=0$. The natural way to treat this “limit circle” is to consider the Čech homology instead of the ordinary singular homology. The Čech homology $\mathrm{\check{H}}_{1}(\mathcal{X})$ is non-trivial, and contains the “limit circle”. In general, a predecessor of the class $\omega$ naturally lives in the Čech homology $\mathrm{\check{H}}_{k}(\overline{\mathcal{N}}_{b_{(\omega,s)}}(\mathcal{X}\subset\mathbf{Y}))$.
To ease the notational burden, we will often drop the parameter $s$ and will use the more succinct notation $b_{\omega}$, $d_{\omega}$, $I_{\omega}$, etc.
Remark 2.20.
A priori, one would expect the collection of all intervals
$\big{\{}I_{\omega};\,\omega\in\mathrm{Spec}_{k}(\Delta_{\bullet})\big{\}}$
to be closely related to the $k$-th persistence diagram of $\Delta_{\bullet}$. Whereas Proposition 2.21 below establishes a sense in which this is the case, it is not always true that all such intervals appear in the interval decomposition of $\operatorname{\tilde{H}}_{k}(\Delta_{\bullet})$ (whenever it exists). An example showing this discrepancy in the case of Vietoris-Rips filtrations can be found in [LMO23, Example 9.2].
Proposition 2.21 ([LMO23, Proposition 9.2]).
Let $\Delta_{\bullet}=\big{(}\Delta_{r},\iota_{r,s}\big{)}_{0<r\leq s}$ be a neighborhood filtration and let $k\geq 1$ be an integer. Then, for all $r<s$, the multiplicity of the interval $(r,s]$ in the barcode of $\mathrm{H}_{k}(\Delta_{\bullet})$ is equal to
$\max\left\{m\in\mathbb{Z}_{\geq 0}\Bigg{|}\begin{array}[]{l}\exists\text{ linearly independent vectors }\omega_{1},\dots,\omega_{m}\in\mathrm{H}_{k}(\Delta_{s})\text{ s.t. }I_{(\omega_{i},s)}=(r,s]\,\forall i\\ \text{ and no nonzero linear combination of these vectors belongs to }\mathrm{Im}((i_{r,s})_{*})\end{array}\right\}.$
This proposition indicates that for each interval $I=(r,s]$ in the barcode of $\Delta_{\bullet}$ there is a finite linearly independent collection $\omega_{1},\ldots,\omega_{m}\in\mathrm{H}_{k}(\Delta_{s})$ satisfying the conditions above such that $I=I_{(\omega_{i},s)}$ for all $i$. One calls any such $\omega_{i}$ a representative of the interval $I$ [^8] In [LMO23, Proposition 9.2] the authors consider the case of the VR filtration of a totally bounded metric space. The same proof applies to the more general statement given above. See [LMO23, page 42] for an example demonstrating the role of the condition that no nonzero linear combination of these vectors belongs to $\mathrm{Im}((\iota_{r,s})_{*})$.
Remark 2.22.
The above proposition implies that if $I$ is an interval in the barcode of a neighborhood filtration, then its length is equal to that of $I_{\omega_{i}}$, where $\omega_{i}$ is as in the statement. Therefore, and as we will do in the rest of the paper, if we have an upper bound for the length of all intervals $I_{\omega}$ where $\omega\in\mathrm{Spec}_{k}(\Delta_{\bullet})$ then we will automatically have an upper bound for the length of every interval in the barcode of $\Delta_{\bullet}$.
Throughout the paper, all homology coefficients are taken as $\mathbb{Z}_{2}$, unless otherwise stated. All results hold for any other coefficient field.
2.2 Geometry of $L^{\infty}(\mathcal{X})$ and Tight Spans
While the discussion in this paper applies to subsets of any Banach space, special attention is paid to the important case of $L^{\infty}$ spaces. There are two main reasons for that. The first one is that if we start with a compact metric space that is a priori not a isometrically embedded into a Banach space, there is a nice way of placing it inside $L^{\infty}(\mathcal{X})$, the space of bounded functions on $\mathcal{X}$ with the supremum norm.
Definition 2.23 (Kuratowski Embedding).
For a compact metric space $(\mathcal{X},\mathbf{d}_{\mathcal{X}})$, the map $\kappa:\mathcal{X}\to L^{\infty}(\mathcal{X})$, defined as $x\mapsto\mathbf{d}_{\mathcal{X}}(x,\cdot)$, is a distance-preserving embedding, and it is called the Kuratowski embedding.
The second reason is that $L^{\infty}$ spaces enjoy the hyperconvexity property: If several balls pairwise intersect then they share a point in common.
Definition 2.24 (Hyperconvex space).
A metric space $E$ is called hyperconvex if for every family $(x_{i},r_{i})_{i\in I}$ of $x_{i}$ in $E$ and $r_{i}\geq 0$ such that $\mathbf{d}_{E}(x_{i},x_{j})\leq r_{i}+r_{j}$ for every $i,j$ in $I$, there exists a point $x\in E$ such that $\mathbf{d}_{E}(x_{i},x)\leq r_{i}$ for every $i$ in $I$.
This implies that the Čech and VR filtrations of a subset of $L^{\infty}(\mathcal{X})$ coincide (up to a factor of two in the filtration index). This gives a way to study VR-lifespans of homology of $\mathcal{X}$ by immersing it in $L^{\infty}(\mathcal{X})$ and then using the tools applicable for Čech-lifespans. Note that in the context of VR-persistence, this idea has been successfully used, cf. 2.32.
The hyperconvexity implies certain universal properties of $L^{\infty}$ spaces, and these properties will be implicitly used below in relation to Urysohn width (see 3.14), and tight spans (a.k.a. injective hulls). We briefly discuss the latter now (for more information we refer the reader to [Lan13]).
Definition 2.25 (Injective metric space).
A metric space $E$ is called injective if for every $1$-Lipschitz map $\phi:\mathcal{X}\to E$ and distance preserving embedding of $\mathcal{X}$ into $\widetilde{\mathcal{X}}$, there exists a $1$-Lipschitz map $\widetilde{\phi}:\widetilde{\mathcal{X}}\to E$ extending $\phi$:
[…]
It turns out that injectivity coincides with hyperconvexity.
Proposition 2.26.
A metric space is injective if and only if it is hyperconvex.
The proof of this proposition can be found in  [AP56] and [Lan13, Proposition 2.3].
It is known that $L^{\infty}(\mathcal{X})$ is injective [Lan13]. However, there exists a more efficient injective space containing $\mathcal{X}$.
Definition 2.27 (Tight span).
The tight span $\operatorname{E}(\mathcal{X})$ of a compact metric space $\mathcal{X}$ is the minimal injective metric space admitting an isometric embedding of $\mathcal{X}$. Minimality here means that any other injective metric space admitting an isometric embedding of $\mathcal{X}$ contains an isometric copy of $\operatorname{E}(\mathcal{X})$.
Tight spans are sometimes called injective hulls or hyperconvex hulls.
Proposition 2.28 (Properties of the tight span [Lan13]).
The tight span $\operatorname{E}(\mathcal{X})$ of a compact metric space $\mathcal{X}$ exists and satisfies:
  1. 1.
    $\operatorname{E}(\mathcal{X})$ is compact.
  2. 2.
    $\operatorname{E}(\mathcal{X})$ is contractible.
  3. 3.
    $\operatorname{E}(\mathcal{X})\cong\mathcal{X}$ for any metric tree $\mathcal{X}$.
  4. 4.
    $\operatorname{diam}(\operatorname{E}(\mathcal{X}))=\operatorname{diam}(\mathcal{X})$.
One particular realization of the tight span $\operatorname{E}(\mathcal{X})$ of $\mathcal{X}$ as a subset of $L^{\infty}(\mathcal{X})$ is given as follows [Lan13, Section 3]):
$\operatorname{E}(\mathcal{X}):=\{f\in\Delta(\mathcal{X}):\text{if }g\in\Delta(\mathcal{X})\text{ and }g\leq f,\text{ then }g=f\,(\text{i.e., }f\text{ is minimal})\}$
where
$\Delta(\mathcal{X}):=\{f\in L^{\infty}(\mathcal{X}):f(x)+f(x^{\prime})\geq\mathbf{d}_{\mathcal{X}}(x,x^{\prime})\text{ for all }x,x^{\prime}\in\mathcal{X}\}.$
Note that from the characterization above, for any $f\in\operatorname{E}(\mathcal{X})$ and any $x\in\mathcal{X}$ it holds that
$f(x)=\max_{x\in\mathcal{X}}\big{(}\mathbf{d}_{\mathcal{X}}(x,x^{\prime})-f(x^{\prime})\big{)}=\big{\|}\mathbf{d}_{\mathcal{X}}(x,\cdot)-f\big{\|}_{\infty}.$ (4)
Remark 2.29.
Not only can the tight span $\operatorname{E}(\mathcal{X})$ be regarded as a subset of $L^{\infty}(\mathcal{X})$ but also, according to [Lan13, Proposition 2.2], there is a $1$-Lipschitz retraction of $L^{\infty}(\mathcal{X})$ to $\operatorname{E}(\mathcal{X})$.
2.3 Filling Radius
The filling radius is a key notion in metric geometry introduced by Gromov as a measure of largeness of a given closed manifold [Gro83]. To fill a manifold $\mathcal{M}$ of dimension $n$, treated as a singular $n$-cycle, means to find an $(n+1)$-dimensional singular chain $\mathcal{D}$ with boundary $\mathcal{M}$, i.e. $\partial\mathcal{D}=\mathcal{M}$. The coefficients of singular homology should be specified, as we discuss below.
Definition 2.30 (Gromov’s filling radius, [Gro83]).
The filling radius $\rho(\mathcal{M})$ of a closed $n$-dimensional Riemannian manifold $\mathcal{M}$ is the infimal number $R>0$ such that $\mathcal{M}$ can be filled inside of the $R$-neighborhood $\mathcal{N}_{R}(\kappa(\mathcal{M}))$ of its Kuratowski image in $L^{\infty}(\mathcal{M})$.
In this definition, the most common convention is to use $\mathbb{Z}$-coefficients if $\mathcal{M}$ is oriented, and $\mathbb{Z}_{2}$-coefficients if $\mathcal{M}$ is not oriented. However, the usual persistence homology is well defined [^9] See, however, [Pat18]. only over fields, so we stick to $\mathbb{Z}_{2}$-coefficients. We note that all homology coefficients in the statements below are $\mathbb{Z}_{2}$, unless otherwise stated.
Before Gromov, a different type of filling radius notion was discussed for submanifolds (or more generally, cycles) in $\mathbb{R}^{N}$ in geometric analysis, especially in the context of isoperimetric problem [BS83, FF60, MS73]. We summarize both types of filling radii in the following definition, adapting it to the context of persistence.
Definition 2.31 (Relative and absolute filling radius of a homology class).
Let $(\mathcal{X},\mathbf{d}_{\mathcal{X}})$ be a metric space, and let $\omega\in\operatorname{\tilde{H}}_{k}(\mathcal{X})$ be a non-trivial reduced singular homology class.
  1. 1.
    Assume additionally that $\mathcal{X}$ is a subset of a Banach space $\mathbf{Y}$, so that the metric $\mathbf{d}_{\mathcal{X}}$ agrees with the one inherited from $\mathbf{Y}$ (that is, the embedding $\mathcal{X}\hookrightarrow\mathbf{Y}$ is distance-preserving). The relative filling radius of $\omega$ is the infimal number $r$ such that the image of $\omega$ under the induced map $\operatorname{\tilde{H}}_{k}(\mathcal{X})\to\operatorname{\tilde{H}}_{k}(\mathcal{N}_{r}(\mathcal{X}))$ is trivial. In other words, it is the infimal size of a neighborhood in $\mathbf{Y}$ where some cycle representing $\omega$ bounds a $(k+1)$-chain. Notation for relative filling radius: $\rho(\omega;\mathcal{X}\subset\mathbf{Y})$.
  2. 2.
    In case when no ambient space $\mathbf{Y}$ is specified, one can take $\mathbf{Y}=L^{\infty}(\mathcal{X})$, and measure the relative filling radius of the Kuratowski image $\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X})$. This way one gets the (absolute) filling radius of $\omega$: $\rho(\omega;\mathcal{X})=\rho(\kappa_{*}\omega;\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))$; cf. [LMO23, Definition 24].
The word “absolute”, which is usually omitted, is justified by the fact that $\rho(\omega;\mathcal{X})$ equals the minimum of $\rho(\iota_{*}\omega;\iota(\mathcal{X})\subset\mathbf{Y})$ over all distance-preserving embeddings $\iota:\mathcal{X}\hookrightarrow\mathbf{Y}$ [Gro83]; we will refer to this as the universal property. Note that when $\mathcal{X}=\mathcal{M}$ is a closed Riemannian manifold, Gromov’s filling radius (2.30) over $\mathbb{Z}_{2}$ is the same as the absolute filling radius of the $\mathbb{Z}_{2}$-fundamental class $[\mathcal{M}]$. Note also that the relationship between the VR filtration of a manifold and its Gromov’s filling radius is studied in [LMO23, Section 9.3].
Proposition 2.32 ([LMO23, Proposition 9.4]).
Let $\mathcal{M}$ be an $n$-dimensional closed connected Riemannian manifold. Then, the interval $I_{\mathcal{M}}:=(0,2\rho(\mathcal{M})]$ appears in the degree-$n$ barcode of $\mathcal{V}_{\bullet}(\mathcal{M})$. Here both the filling radius and persistent homology are computed with $\mathbb{Z}_{2}$ coefficients.
Remark 2.33 (Relative Filling Radius vs. Absolute Filling Radius).
To give an idea about the difference between relative filling radius in $\mathbf{Y}$ and absolute filling radius, here we give a toy example of a flying saucer in $\mathbf{Y}=\mathbb{R}^{3}$ (with Euclidean metric). Let $\mathrm{E}$ be the ellipsoid in $\mathbb{R}^{3}$ given by $\mathrm{E}=\{(x,y,z)\mid\frac{x^{2}}{100}+\frac{y^{2}}{100}+\frac{z^{2}}{1}=1\}$, and let $[\mathrm{E}]$ be its fundamental class in $\mathrm{H}_{2}(\mathrm{E})$. There are two different ways to treat $\mathrm{E}$ as a metric space, resulting in the different values of its filling radius.
  1. 1.
    One way is to consider the intrinsic Riemannian metric $g$ on $\mathrm{E}$ induced by the Euclidean metric of $\mathbb{R}^{3}$. Thus obtained Riemannian surface $\mathrm{E}_{g}$ has the absolute filling radius about $10$: $\rho([\mathrm{E}];\mathrm{E}_{g})\sim 10$ (the exact computation is tricky).
  2. 2.
    The other way is to borrow the extrinsic distance function from $\mathbb{R}^{3}$. This does not make $\mathrm{E}$ a Riemannian manifold, but rather just a compact metric space embedded in $\mathbb{R}^{3}$ in a distance-preserving way. The corresponding relative filling radius equals $1$: $\rho([\mathrm{E}];\mathrm{E}\subset\mathbb{R}^{3})=1$.
In this example we have $\rho([\mathrm{E}];\mathrm{E}_{g})>\rho([\mathrm{E}];\mathrm{E}\subset\mathbb{R}^{3})$, which might seem to contradict the note above saying that $\rho(\omega;\mathcal{X})\leq\rho(\omega;\mathcal{X}\subset\mathbf{Y})$ for distance-preserving embeddings $\mathcal{X}\subset\mathbf{Y}$. There is no contradiction here: even though the embedding $\mathrm{E}_{g}\subset\mathbb{R}^{3}$ is a Riemannian isometry, it is not distance-preserving. Indeed, the extrinsic distance in $\mathbb{R}^{3}$ between two points of $\mathrm{E}$ is smaller than the intrinsic distance between them inside $\mathrm{E}$ (the length of the shortest path in $\mathrm{E}$). If one computes the filling radius of $E$ with the extrinsic metric of $\mathbb{R}^{3}$, it will be at most $1$.
In the following sections, we will see that Gromov’s filling radius of a Riemannian manifold isometrically embedded in an ambient space $\mathbf{Y}$ can highly overestimate the lifespan of a topological feature in the Čech filtration and that the relative filling radius is better adapted to this context.
Filling Radius & Persistence. The notion of lifespan of a homology class (2.18) is directly related to the filling radius as follows. Let $\mathcal{X}$ be a compact subset of a Banach space $(\mathbf{Y},\|\cdot\|)$. Let $(b_{\omega},d_{\omega})$ be the homological birth and death times of a degree-$k$ homology class $\omega$ present at time $s$ in the neighborhood filtration $\mathcal{N}_{\bullet}(\mathcal{X}\subset\mathbf{Y})$. Right from the definition of the relative filling radius, we get
$d_{\omega}-s=\rho(\omega;\mathcal{N}_{s}(\mathcal{X}\subset\mathbf{Y})\subset\mathbf{Y}).$
A bit more generally, if $\omega_{r}\in\operatorname{\tilde{H}}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y}))$ is a predecessor of $\omega$ (that is, it is mapped to $\omega$ by the map in homology induced by the inclusion), then
$d_{\omega}-r=\rho(\omega_{r};\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})\subset\mathbf{Y}).$
Letting $r\to b_{\omega}$, we get a formula for the lifespan of $\omega$ in terms of the filling radii of the predecessors of $\omega$:
$d_{\omega}-b_{\omega}$ $=\sup_{b_{\omega}<r\leq s}\rho(\omega_{r};\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})\subset\mathbf{Y})$
$=\lim_{r\to b_{\omega}+0}\rho(\omega_{r};\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})\subset\mathbf{Y}).$ ($\star$)
As clarified in 2.19, this lifespan cannot be written in term of the filling radius of homology of $\overline{\mathcal{N}}_{b_{\omega}}(\mathcal{X}\subset\mathbf{Y})$. We cannot expect (a predecessor of) $\omega$ to be present in the homology of $\overline{\mathcal{N}}_{b_{\omega}}(\mathcal{X}\subset\mathbf{Y})$, unless we work with Čech homology, and redefine the filling radius correspondingly.
Formula (2.3) was given for the neighborhood filtration. In view of the functorial nerve theorem (2.6, 2.7), the formula specializes to two important cases:
  • For any $\omega\in\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$,
    $d_{\omega}-b_{\omega}=\lim_{r\to b_{\omega}+0}\rho(\omega_{r};\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})\subset\mathbf{Y}).$
  • For any $\omega\in\mathrm{Spec}_{k}(\mathcal{V}_{\bullet}(\mathcal{X}))$,
    $d_{\omega}-b_{\omega}=2\lim_{r\to b_{\omega}+0}\rho(\omega_{r};\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))\subset L^{\infty}(\mathcal{X})).$
Remark 2.34.
For the particular case of VR-lifespans, i.e. when $\mathbf{Y}=L^{\infty}(\mathcal{X})$, it is important to understand the behavior or the filling radii $\rho(\omega_{r};\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))\subset L^{\infty}(\mathcal{X}))$. We make a comment that this cumbersome notation is a bit redundant; it turns out that
$\rho(\omega_{r};\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))\subset L^{\infty}(\mathcal{X}))=\rho(\omega_{r};\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))).$
This is not obvious by default, since the absolute filling radius in the right-hand side should be computed in the space $L^{\infty}(\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X})))$. Nevertheless, this equality holds true. The inequality $\geq$” follows from the universal property. The inequality $\leq$” can be easily shown by making use of the evident $1$-Lipschitz restriction map $L^{\infty}(\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X})))\to L^{\infty}(\mathcal{X})$.
Auxiliary definitions. Before finishing the background section, we define two versions of the radius of a set $\mathcal{X}$, which will be used in the remainder of the paper. In the first one, there is no reference to an ambient space, and the center is in the set $\mathcal{X}$. In the second one (circumradius), the radius of $\mathcal{X}$ is computed in an ambient space, and the center may not be in $\mathcal{X}$.
Definition 2.35 (Radius).
Let $(\mathcal{X},\mathbf{d}_{\mathcal{X}})$ be a compact metric space.
  1. 1.
    The radius of $\mathcal{X}$ is
    $\mathrm{rad}(\mathcal{X})=\inf_{y\in\mathcal{X}}\sup_{x\in\mathcal{X}}\mathbf{d}_{\mathcal{X}}(x,y).$
  2. 2.
    Assume additionally that $\mathcal{X}$ is a compact subset of a Banach space $\mathbf{Y}$, so that the metric $\mathbf{d}_{\mathcal{X}}$ agrees with the one inherited from $\mathbf{Y}$ (that is, the embedding $\mathcal{X}\hookrightarrow\mathbf{Y}$ is distance-preserving). The circumradius of $\mathcal{X}$ in $\mathbf{Y}$ is
    $\mathrm{rad}(\mathcal{X}\subset\mathbf{Y})=\inf_{y\in\mathbf{Y}}\sup_{x\in\mathcal{X}}\|x-y\|.$
3 Bounding Lifespans via Widths
In this section, we recall and establish upper bounds the filling radius, which can be used to estimate the lifespans in both Čech and Vietoris–Rips settings. The basic idea behind many constructions is simple and can be illustrated as follows. Let $\mathcal{X}$ be a subset of a Banach space $\mathbf{Y}$, and let $k>\ell$ be positive integers. A degree-$k$ homology class of $\mathcal{X}$ can be “killed” by deforming $\mathcal{X}$ to an $\ell$-dimensional complex inside $\mathbf{Y}$ (a “core”), and if every point moves by some controlled distance, then we get an estimate for the filling radius of the degree-$k$ homology.
3.1 Background on Widths
Here we recall three classical approaches to measuring approximate dimension. Informally, the $k$-width of a space $\mathcal{X}$ measures to which extent $\mathcal{X}$ fails to be $k$-dimensional.
The Urysohn width was historically first. The following definition is equivalent to the one given by Pavel Urysohn around 1923 [Ale26] in the context of dimension theory.
Definition 3.1 (Urysohn width).
Let $\mathcal{X}$ be a compact metric space. For an integer $k\geq 0$, the Urysohn $k$-width of $\mathcal{X}$ is defined as
$\operatorname{UW}_{k}(\mathcal{X})=\inf_{f}\sup_{p}\operatorname{diam}(f^{-1}(p)),$
where $f:\mathcal{X}\to\Delta^{k}$ is any continuous map to any finite $k$-dimensional simplicial complex.
By definition, widths are monotone in the dimension parameter:
$\operatorname{UW}_{0}(\mathcal{X})\geq\operatorname{UW}_{1}(\mathcal{X})\geq\operatorname{UW}_{2}(\mathcal{X})\geq\dots,$
and $\operatorname{UW}_{0}(\mathcal{X})=\operatorname{diam}(\mathcal{X})$ if $\mathcal{X}$ is connected. The $n$-width of an $n$-dimensional manifold is zero, and all preceding $k$-widths are positive for $0\leq k<n$.
The crucial connection between the Urysohn width and the filling radius is due to Gromov.
Theorem 3.2 ([Gro83]).
For any closed $n$-dimensional Riemannian manifold $\mathcal{M}$,
$\rho(\mathcal{M})\leq\frac{1}{2}\operatorname{UW}_{n-1}(\mathcal{M}).$
The coefficients in the definition of $\rho(\mathcal{M})$ do not matter here; everywhere below we stick to the case of $\mathbb{Z}_{2}$-coefficients since this is the main case of interest for persistent homology. The importance of this result for bounding lifespans becomes immediate once one notices that the proof of this inequality can be easily generalized for any topological feature, and not just the fundamental class.
Theorem 3.3.
For any compact metric space $\mathcal{X}$, any integer $k\geq 1$, and any homology class $\omega\in\operatorname{\tilde{H}}_{k}(\mathcal{X})$,
$\rho(\omega;\mathcal{X})\leq\frac{1}{2}\operatorname{UW}_{k-1}(\mathcal{X}).$
Proof sketch, following [Gro20].
Let $f:\mathcal{X}\to\Delta^{k-1}$ be a map for which $\delta=\sup_{p}\operatorname{diam}(f^{-1}(p))$ is just a tiny amount bigger than $\operatorname{UW}_{k-1}(\mathcal{X})$. Consider the cylinder $\mathcal{X}\times[0,\delta/2]$ and glue its end $\mathcal{X}\times\{\delta/2\}$ to a copy of $\Delta^{k-1}$ along the map $f$; that means, pinch every fiber of $f$ inside $\mathcal{X}\times\{\delta/2\}$ to a point. The resulting space $\mathcal{C}_{f}$ can be endowed with a metric that restricts on $\mathcal{X}\times\{0\}$ to the original metric of $\mathcal{X}$, and makes the length of every interval $\{x\}\times[0,\delta/2]$ equal to $\delta/2$. Now, any $k$-cycle $S$ in $\mathcal{X}\times\{0\}$ representing $\omega$ becomes null-homologous in $\mathcal{C}_{f}$ (informally, one can just slide it towards the pinched end of the cylinder, where it degenerates to a $(k-1)$-dimensional set). Next, we embed $\mathcal{C}_{f}$ to $L^{\infty}(\mathcal{C}_{f})$ in a distance-preserving way, and post-compose it with the $1$-Lipschitz restriction $L^{\infty}(\mathcal{C}_{f})\to L^{\infty}(\mathcal{X})$, corresponding to the inclusion $\mathcal{X}\times\{0\}\subset\mathcal{C}_{f}$. The $(k+1)$-chain that we built in $\mathcal{C}_{f}$ to fill $S$ pushes forward to $L^{\infty}(\mathcal{X})$, and there it lies within distance $\delta/2$ of the Kuratowski image of $\mathcal{X}$. ∎
It is natural to look for a relative version of Theorem 3.3. We provide such an inequality below (Theorem 3.5), additionally replacing its right-hand side by the quantity that is comparable to the Urysohn width but easier to compute; the corresponding metric invariant implicitly appeared in the work of Pavel Alexandrov [Ale33] on dimension theory.
Definition 3.4 (Alexandrov width).
Let $\mathcal{X}$ be a compact subset of a Banach space $(\mathbf{Y},\|\cdot\|)$. For an integer $k\geq 0$, the Alexandrov $k$-width of $\mathcal{X}$ (relative to $\mathbf{Y}$) is defined as
$\operatorname{AW}_{k}(\mathcal{X}\subset\mathbf{Y})=\inf_{f}\sup_{x\in\mathcal{X}}\|x-f(x)\|,$
where $f:\mathcal{X}\to\mathbf{Y}$ is any continuous map whose image is a finite simplicial complex of dimension at most $k$.
The Alexandrov width enjoys similar properties with the Urysohn width: it is monotonically decreasing in $k$ until it reaches zero when $k$ becomes equal the dimension of $\mathcal{X}$. If $\mathcal{X}$ is connected, $\operatorname{AW}_{0}(\mathcal{X}\subset\mathbf{Y})$ equals the circumradius of $\mathcal{X}$ in $\mathbf{Y}$, that is, $\operatorname{AW}_{0}(\mathcal{X}\subset\mathbf{Y})=\mathrm{rad}(\mathcal{X}\subset\mathbf{Y})$ (see 2.35).
It is easy to see directly from the definitions that $\operatorname{UW}_{k}(\mathcal{X})\leq 2\operatorname{AW}_{k}(\mathcal{X}\subset\mathbf{Y})$. Combining this with Theorem 3.3, we get the estimate $\rho(\omega;\mathcal{X})\leq\operatorname{AW}_{k-1}(\mathcal{X}\subset\mathbf{Y})$, which we will improve by replacing the absolute filling radius by the relative one.
Theorem 3.5.
For any compact set $\mathcal{X}$ sitting in a Banach space $\mathbf{Y}$, any integer $k\geq 1$, and any homology class $\omega\in\mathrm{H}_{k}(\mathcal{X})$,
$\rho(\omega;\mathcal{X}\subset\mathbf{Y})\leq\operatorname{AW}_{k-1}(\mathcal{X}\subset\mathbf{Y}).$
This is the relative analogue of Theorem 3.3. It will follow from a stronger estimate below (Theorem 3.10), but the intuition behind it is simple as described at the preamble of this section: One can kill higher homology of $\mathcal{X}$ by deforming it to a low-dimensional complex in $\mathbf{Y}$, and if every point moves by some controlled distance, then we get an estimate for the filling radius.
The virtues of 3.5 are twofold. First, the Alexandrov width seems to be easier to estimate than the Urysohn width. Basically, the Urysohn width considers all maps to something $k$-dimensional, whereas the Alexandrov width only considers a special class of those maps with the images lying in the same ambient space $\mathbf{Y}$; this is also the reason why $\operatorname{UW}_{k}(\mathcal{X})\leq 2\operatorname{AW}_{k}(\mathcal{X}\subset\mathbf{Y})$. Second, in Appendix A we explain that $\operatorname{AW}_{k}(\mathcal{X}\subset\mathbf{Y})\leq\operatorname{UW}_{k}(\mathcal{X})$. Hence, with 3.5, this implies one can bound the relative filling radius with the Urysohn width (if $\mathcal{X}$ is compact), too, and we do not lose much when substituting the widths, since
$\operatorname{AW}_{k}(\mathcal{X}\subset\mathbf{Y})\leq\operatorname{UW}_{k}(\mathcal{X})\leq 2\operatorname{AW}_{k}(\mathcal{X}\subset\mathbf{Y}).$
When $\mathbf{Y}$ is hyperconvex (for example, $L^{\infty}(\mathcal{X})$), one can claim more (see A.2):
$\operatorname{UW}_{k}(\mathcal{X})=2\operatorname{AW}_{k}(\mathcal{X}\subset\mathbf{Y}).$
While Urysohn’s and Alexandrov’s notions measure non-linear width, a simpler concept of linear width appeared in the work of Kolmogorov [Kol36] in the context of approximation theory. It is commonly used in infinite-dimensional settings.
Definition 3.6 (Kolmogorov width).
Let $\mathcal{X}$ be a compact subset of a Banach space $(\mathbf{Y},\|\cdot\|)$. For an integer $k\geq 0$, the Kolmogorov $k$-width of $\mathcal{X}$ (relative to $\mathbf{Y}$) is defined as
$\operatorname{KW}_{k}(\mathcal{X}\subset\mathbf{Y})=\inf\{r~{}|~{}\mathcal{X}\subset\mathcal{N}_{r}(\mathcal{P}\subset\mathbf{Y})\text{ for some affine $k$-plane }\mathcal{P}\subset\mathbf{Y}\}.$
Again, these quantities monotonically decrease in $k$. It is immediate from definitions that $\operatorname{AW}_{k}(\mathcal{X}\subset\mathbf{Y})\leq\operatorname{KW}_{k}(\mathcal{X}\subset\mathbf{Y})$. From Theorem 3.5 it follows then the relative filling radius of degree-$k$ homology can be bounded as $\rho(\cdot;\mathcal{X}\subset\mathbf{Y})\leq\operatorname{KW}_{k-1}(\mathcal{X}\subset\mathbf{Y})$. The following finer estimate is another corollary of Theorem 3.10.
Theorem 3.7.
Let $\mathcal{X}$ be a compact subset of a Banach space $\mathbf{Y}$. Then for any homology class $\omega\in\operatorname{\tilde{H}}_{k}(\mathcal{X})$, $k\geq 0$,
$\rho(\omega;\mathcal{X}\subset\mathbf{Y})\leq\operatorname{KW}_{k}(\mathcal{X}\subset\mathbf{Y}).$
Remark 3.8 (Filling radius and $k$-widths).
We give a toy example to illustrate the relation between the $k$-width and filling radius.
For $m\leq N$, consider the $(m-1)$-dimensional ellipsoid $\mathrm{E}\subset\mathbb{R}^{N}$ (with the extrinsic metric) given by
$\mathrm{E}=\left\{\mathbf{x}\in\mathbb{R}^{N}~{}\middle|~{}\sum_{i=1}^{m}\dfrac{x_{i}^{2}}{a_{i}^{2}}=1,\mbox{ and }x_{j}=0\mbox{ for }j>m\right\},$
where $a_{1}>a_{2}>\dots>a_{m}>0$. The widths can be roughly estimated as $\operatorname{UW}_{k}(\mathrm{E})\sim\operatorname{AW}_{k}(\mathrm{E}\subset\mathbb{R}^{N})\sim a_{k+1}$ for $0\leq k<m$. On the other hand, the relative filling radius of $\mathrm{E}$ in $\mathbb{R}^{N}$ equals the length of the shortest axis: $\rho([\mathrm{E}];\mathrm{E}\subset\mathbb{R}^{N})=a_{m}$. Here, the dimensions $m,N$ and the sequence $\{a_{i}\}$ are arbitrary. 3.5 tells us that $\rho([\mathrm{E}];\mathrm{E}\subset\mathbb{R}^{N})\leq\operatorname{AW}_{m-2}(\mathrm{E}\subset\mathbb{R}^{N})$, but we see that the discrepancy between $\rho([\mathrm{E}];\mathrm{E}\subset\mathbb{R}^{N})=a_{m}$ and $\operatorname{AW}_{m-2}(\mathrm{E}\subset\mathbb{R}^{N})\sim a_{m-1}$ can be arbitrarily large. In other words, the Alexandrov (and Urysohn) width can highly overestimate the filling radius. However, the estimate of 3.7 in this example is sharp: $\rho([\mathrm{E}];\mathrm{E}\subset\mathbb{R}^{N})=\operatorname{KW}_{m-1}(\mathrm{E}\subset\mathbb{R}^{N})=a_{m}$
For more information on widths, we refer the reader to [Bal21].
3.2 Bounds for Čech Lifespans via Treewidth
Recall (from formula (2.3) in Section 2.3) that $\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$-lifespans can be written in terms of relative filling radii as follows:
$d_{\omega}-b_{\omega}=\lim_{r\to b_{\omega}+0}\rho(\omega_{r};\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})\subset\mathbf{Y}),$
where $\omega_{r}$ maps to $\omega$ by the homology map induced by the inclusion (see Section 2.3). To estimate lifespans, we need upper bounds for filling radii, such as in 3.5 and 3.7. Both estimates can be simultaneously strengthened using the following new width invariant. Recall that throughout the paper, we use homology with $\mathbb{Z}_{2}$-coefficients (though the same definition makes sense with any coefficients).
Definition 3.9 (Treewidth).
[…]
$\operatorname{TW}_{k}(\mathcal{X}\subset\mathbf{Y})=\inf_{f}\sup_{x\in\mathcal{X}}\|x-f(x)\|,$
[…]
The treewidth satisfies evident properties:
$\mathrm{rad}(\mathcal{X}\subset\mathbf{Y})=\operatorname{TW}_{0}(\mathcal{X}\subset\mathbf{Y})\geq\operatorname{TW}_{1}(\mathcal{X}\subset\mathbf{Y})\geq\operatorname{TW}_{2}(\mathcal{X}\subset\mathbf{Y})\geq\ldots.$
[…]
The motivation behind the definition of treewidth is to give a common generalization of 3.5 and 3.7. Both follow from the following result combined with the evident bounds $\operatorname{TW}_{k}(\mathcal{X}\subset\mathbf{Y})\leq\operatorname{AW}_{k-1}(\mathcal{X}\subset\mathbf{Y})$ and $\operatorname{TW}_{k}(\mathcal{X}\subset\mathbf{Y})\leq\operatorname{KW}_{k}(\mathcal{X}\subset\mathbf{Y})$.
Theorem 3.10.
For any compact set $\mathcal{X}$ sitting in a Banach space $\mathbf{Y}$, any integer $k\geq 0$, and any homology class $\omega\in\operatorname{\tilde{H}}_{k}(\mathcal{X})$, $k\geq 0$,
$\rho(\omega;\mathcal{X}\subset\mathbf{Y})\leq\operatorname{TW}_{k}(\mathcal{X}\subset\mathbf{Y}).$
Proof.
Fix any number $\delta>\operatorname{TW}_{k}(\mathcal{X}\subset\mathbf{Y})$ and pick a witness map $f:\mathcal{X}\to\mathbf{Y}$ with at most $k$-dimensional image $\mathcal{T}$, such that $\|x-f(x)\|\leq\delta$ for all $x\in\mathcal{X}$. Pick a cycle $S\subset\mathcal{X}$ representing $\omega$ and continuously deform it via the linear homotopy $h:S\times[0,1]\to\mathbf{Y}$ given by
$h(x,t)=(1-t)x+tf(x).$
Every point $x\in S$ moves by distance at most $\delta$, so the continuous deformation always stays in $\mathcal{N}_{\delta}(S\subset\mathbf{Y})$. At time $t=1$, the deformed cycle $h(S,1)$ lies in $\mathcal{T}$ and bounds a $(k+1)$ chain within $h(S,1)$, because $\mathcal{T}$ is of dimension $k$ with trivial $\operatorname{\tilde{H}}_{k}(\mathcal{T})$. Therefore, $S$ is null-homologous in $\mathcal{N}_{\delta}(S\subset\mathbf{Y})$ (basically, $S$ bounds the $(k+1)$-dimensional trace of the homotopy $h$). ∎
A simple example in which 3.10 gives a stronger bound than 3.5 and 3.7 is depicted in Figure 1; the fundamental class of this circle has a small filling radius, which can be efficiently estimated by the $1$-treewidth (approximating the shape by a tree, red in the figure), while approximations by straight lines or points give significantly worse bounds.
Figure 1: A shape with small $1$-treewidth but considerable Alexandrov 00-width and Kolmogorov $1$-width
We proceed with a strengthening applicable to non-compact sets of the form $\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})$. Notice that the width in the right-hand side of the following estimate is computed on a compact set, since in Banach spaces the closure of the convex hull of a compact set is compact [Lax02, Section 13, Exercise 9].
Theorem 3.11.
For any compact set $\mathcal{X}$ sitting in a Banach space $\mathbf{Y}$, any integer $k\geq 0$, any non-negative number $r$, and any homology class $\omega\in\operatorname{\tilde{H}}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y}))$,
$\rho(\omega;\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})\subset\mathbf{Y})\leq\operatorname{TW}_{k}(\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\overline{\mathrm{conv}}(\mathcal{X}))\subset\mathbf{Y}).$
Proof.
Let our homology class $\omega\in\operatorname{\tilde{H}}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y}))$ be represented by a singular cycle $S_{0}$ supported in $\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})$. Since $S_{0}$ is a compact subset of $\mathbf{Y}$, it can be covered by finitely many open balls $B_{r}(x),x\in\mathcal{X}^{\prime}\subset\mathcal{X},|\mathcal{X}^{\prime}|<\infty$ (here $B_{r}(x)$ denotes the open ball of radius $r$ centered at $x$). Consider the union $U=\bigcup\limits_{x\in\mathcal{X}^{\prime}}B_{r}(x)$ and choose any partition of unity subordinate to the cover of $U$ by $B_{r}(x),x\in\mathcal{X}^{\prime}$. This partition of unity gives rise to a map from $U$ to the nerve $Z$ of that cover. By the nerve theorem this map $\phi:U\to Z$ is a homotopy equivalence, with an evident homotopy inverse $\psi:Z\to U$, which sends every vertex of $Z$ to the center of the corresponding ball in the cover $U=\bigcup\limits_{x\in\mathcal{X}^{\prime}}B_{r}(x)$, and extends affinely on the rest of $Z$. Therefore, $\omega|_{U}=\psi_{*}\phi_{*}\omega|_{U}\in\operatorname{\tilde{H}}_{k}(U)$, and the class $\omega$ can be represented by a cycle $S$ supported in $\psi(Z)\subset\overline{\mathrm{conv}}(\mathcal{X})$. A similar argument shows that is $S$ bounds in $\mathcal{N}_{R}(\mathcal{X}\subset\mathbf{Y})$, for some $R>r$, then the filling chain can be also taken with the support in $\overline{\mathrm{conv}}(\mathcal{X})$. Therefore,
$\rho(\omega;\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})\subset\mathbf{Y})\leq\rho(\omega;\mathcal{N}_{r}(\mathcal{X}\subset\overline{\mathrm{conv}}(\mathcal{X}))\subset\mathbf{Y}).$
Now we are in position to apply 3.10 to the compact set $\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\overline{\mathrm{conv}}(\mathcal{X}))$:
$\rho(\omega;\mathcal{N}_{r}(\mathcal{X}\subset\overline{\mathrm{conv}}(\mathcal{X}))\subset\mathbf{Y})$ $=\rho(\omega;\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\overline{\mathrm{conv}}(\mathcal{X}))\subset\mathbf{Y})$
$\leq\operatorname{TW}_{k}(\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\overline{\mathrm{conv}}(\mathcal{X}))\subset\mathbf{Y}).$
Corollary 3.12 (Čech Lifespans via Treewidth).
Let $\mathcal{X}$ be compact subset of a Banach space $\mathbf{Y}$ and let $\omega\in\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$, $k\geq 0$. Then,
$d_{\omega}-b_{\omega}\leq\operatorname{TW}_{k}(\overline{\mathcal{N}}_{b_{\omega}}(\mathcal{X}\subset\overline{\mathrm{conv}}(\mathcal{X}))\subset\mathbf{Y}).$
In particular, $d_{\omega}-b_{\omega}\leq\operatorname{TW}_{k}(\overline{\mathrm{conv}}(\mathcal{X})\subset\mathbf{Y})$.
Remark 3.13.
Recall the monotonicity of the treewidth: $\operatorname{TW}_{k}(\cdot)\leq\operatorname{TW}_{0}(\cdot)=\mathrm{rad}(\cdot)$. Therefore, as a trivial consequence of 3.12, we can upper-bound lifespans in all dimensions by $\operatorname{TW}_{0}(\overline{\mathrm{conv}}(\mathcal{X})\subset\mathbf{Y})=\mathrm{rad}(\mathcal{X}\subset\mathbf{Y})$.
Proof of 3.12.
We use the aforementioned formula
$d_{\omega}-b_{\omega}=\lim_{r\to b_{\omega}+0}\rho(\omega_{r};\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})\subset\mathbf{Y}),$
and bound each filling radius by the treewidth of $\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})$:
$\rho(\omega_{r};\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})\subset\mathbf{Y})\leq\operatorname{TW}_{k}(\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\overline{\mathrm{conv}}(\mathcal{X}))\subset\mathbf{Y}).$
To conclude, we need a continuity property:
$\lim_{r\to b_{\omega}+0}\operatorname{TW}_{k}(\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\overline{\mathrm{conv}}(\mathcal{X}))\subset\mathbf{Y})=\operatorname{TW}_{k}(\overline{\mathcal{N}}_{b_{\omega}}(\mathcal{X}\subset\overline{\mathrm{conv}}(\mathcal{X}))\subset\mathbf{Y}).$
Its proof is explained in the appendix, see A.3. ∎
Combining this with the trivial inequality $\operatorname{TW}_{k}(\cdot)\leq\operatorname{AW}_{k-1}(\cdot)$, we obtain the following bound.
[…]
Figure 2: For the surface $S$ above, while $\operatorname{AW}_{1}(S\subset\mathbb{R}^{3})$ is small, the lifespan of red curve $\alpha\in\operatorname{\tilde{H}}_{1}(S)$ is large.
We would like to point out that while $k$-widths are effective for bounding lifespans of homology classes of degree $>k$, they do not say much about the lifespans of homology classes in lower degrees. In Figure 2, we give a simple $2$-dimensional example $S$. Here, $\operatorname{AW}_{1}(S\subset\mathbb{R}^{3})$ is small, and hence, the homology class in $\operatorname{\tilde{H}}_{2}(S)$ has short lifespans by Section 1. However, one can easily see that homology classes $\alpha\in\operatorname{\tilde{H}}_{1}(S)$ can still have very large lifespans. One can generalize this example to any dimension and codimension. In Section 5 we take a different perspective, and give global bounds for lifespans and death times in all degrees simultaneously.
3.3 Bounds for Vietoris–Rips Lifespans via Urysohn Width
Recall the relation between $\mathrm{Spec}_{k}(\mathcal{V}_{\bullet}(\mathcal{X}))$-lifespans and filling radii implied by formula (2.3) of Section 2.3:
$d_{\omega}-b_{\omega}=2\lim_{r\to b_{\omega}+0}\rho(\omega_{r};\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))\subset L^{\infty}(\mathcal{X})),$
where $\omega_{r}$ maps to $\omega$ by the homology level map induced by the inclusion, and $\kappa:\mathcal{X}\to L^{\infty}(\mathcal{X})$ is the Kuratowski embedding. Recall also (2.34) that
$\rho(\omega_{r};\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))\subset L^{\infty}(\mathcal{X}))=\rho(\cdot;\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))).$
To estimate VR-lifespans, we need to know how to upper-bound quantities like $\rho(\cdot;\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X})))$, for which Theorem 3.3 is not applicable: the set $\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))$ is not compact (even after taking the closure). One way to deal with non-compactness is to intersect $\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X}))$ with the closure of the convex hull of $\mathcal{X}$, like we did in 3.11. That theorem, together with the inequality $\operatorname{TW}_{k}(\cdot)\leq\operatorname{AW}_{k-1}(\cdot)$, will give us a bound on $\rho(\cdot;\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X})))$ in terms of the Alexandrov width of $\overline{\mathcal{N}}_{r}(\kappa(\mathcal{X})\subset\overline{\mathrm{conv}}(\kappa(\mathcal{X})))$. It turns out the latter can be interpreted in terms of the Urysohn width.
Remark 3.14 ($L^{\infty}$-interpretation of Urysohn width).
In the inequality $\rho(\omega;\mathcal{X})\leq\frac{1}{2}\operatorname{UW}_{k-1}(\mathcal{X})$, the left-hand side is defined via the Kuratowski embedding in $L^{\infty}(\mathcal{X})$, and it is instructive to interpret the right-hand side as well, in a way compatible with the picture of $\kappa(\mathcal{X})$ sitting in $L^{\infty}(\mathcal{X})$. This interpretation goes back to V. Tikhomirov [Tik76]. Some references in English are [Gro83, Appendix 1, Proposition (D)] or [Bal21, Theorem 2.1.9 and Definition 2.3.1]; we provide here a short summary. The width $\operatorname{UW}_{k-1}(\mathcal{X})$ can be equivalently defined as the infimal number $\delta>0$ such that there is map $f:\mathcal{X}\to L^{\infty}(\mathcal{X})$ whose image is at most $(k-1)$-dimensional complex in $L^{\infty}(\mathcal{X})$, and such that $\|\kappa(x)-f(x)\|_{\infty}\leq\delta/2$ for all $x\in\mathcal{X}$. In a nutshell, Theorem 3.3 says that one can try killing higher homology of $\mathcal{X}$ by deforming $\kappa(\mathcal{X})$ to a low-dimensional complex inside $L^{\infty}(\mathcal{X})$, and if every point moves by some controlled distance, then we get an estimate for the filling radius. This interpretation can be shortly reformulated as follows:
$\operatorname{UW}_{k-1}(\mathcal{X})=2\operatorname{AW}_{k-1}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X})).$
Given that interpretation, the following estimate is immediate:
$\rho(\omega;\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X})))\leq\frac{1}{2}\operatorname{UW}_{k-1}(\overline{\mathcal{N}}_{r}(\kappa(\mathcal{X})\subset\overline{\mathrm{conv}}(\kappa(\mathcal{X})))).$
We will go further and estimate $\rho(\cdot;\mathcal{N}_{r}(\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X})))$ even better, by replacing the convex hull with the tight span (2.27). The following result subsumes Theorem 3.3 when $r=0$. The specific realization of $\operatorname{E}(\mathcal{X})$ in $L^{\infty}(\mathcal{X})$ that was described in Section 2.2 has an isometric copy of $\mathcal{X}$ inside, and the corresponding embedding of $\mathcal{X}$ in $L^{\infty}(\mathcal{X})$ agrees with the Kuratowski embedding. To simplify notation a bit, we omit $\kappa$ and simply write $\mathcal{X}\subset\operatorname{E}(\mathcal{X})\subset L^{\infty}(\mathcal{X})$ in the rest of this section. Notice that the set $\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))$ is compact, as a closed subset of the compact metric space $\operatorname{E}(\mathcal{X})$.
Theorem 3.15.
For any compact metric space $\mathcal{X}$, any integer $k\geq 1$, any real number $r\geq 0$, and any homology class $\omega\in\mathrm{H}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X})))$,
$\rho(\omega;\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X})))\leq\frac{1}{2}\operatorname{UW}_{k-1}(\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))).$
Proof.
We have a chain of distance-preserving embeddings $\mathcal{X}\subset\operatorname{E}(\mathcal{X})\subset L^{\infty}(\mathcal{X})$ as well as $\mathcal{X}\subset\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\subset\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))$. All distances below are measured using the $L^{\infty}$-norm. We need to bound $\rho(\omega;\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))\subset L^{\infty}(\mathcal{X}))$.
Recall the following property of the tight span: there exists a $1$-Lipschitz retraction $\pi:L^{\infty}(\mathcal{X})\to\operatorname{E}(\mathcal{X})$. Our first observation is that $\pi$ maps $\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))$ to $\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))$. Indeed, take any point $y\in\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))$. Since $\|y-x\|<r$ for some $x\in\mathcal{X}$, and $\pi$ is $1$-Lipschitz retraction, it follows that $\|\pi(y)-x\|=\|\pi(y)-\pi(x)\|\leq\|y-x\|<r$. Therefore, $\pi(\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X})))=\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))$. Denote the reverse inclusion $\iota:\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\to\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))$. Since both $y$ and $\pi(y)$ lie in the open radius $r$ ball centered at $x$, the whole straight line segment between them lies in that ball and in $\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))$. Therefore, a cycle representing $\omega$ can be continuously deformed to a cycle representing $\iota_{*}\pi_{*}\omega$ by letting each $y$ in the support of $\omega$ slide along the straight line segment towards $\pi(y)$. This homotopy happens entirely in $\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))$, and therefore, $\omega=\iota_{*}\pi_{*}\omega$.
It should be obvious now that
$\rho(\omega;\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))\subset L^{\infty}(\mathcal{X}))\leq\rho(\pi_{*}\omega;\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\subset L^{\infty}(\mathcal{X})).$
Indeed, any cycle $S$ representing $\pi_{*}\omega\in\operatorname{\tilde{H}}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X})))$ lies entirely in $\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))$ and represents there the homology class $\iota_{*}\pi_{*}\omega=\omega\in\operatorname{\tilde{H}}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X})))$. If $S$ bounds in the neighborhood of $\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))$ of certain radius, then it bounds even in a smaller neighborhood of $\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))$.
Using the evident $1$-Lipschitz restriction map $L^{\infty}(\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X})))\to L^{\infty}(\mathcal{X})$, it is easy to show that
$\rho(\pi_{*}\omega;\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\subset L^{\infty}(\mathcal{X}))\leq\rho(\pi_{*}\omega;\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))).$
(In fact, the equality holds here, but we only need the inequality.) To show this inequality, we can just use that $1$-Lipschitz map to send cycles and filling chains in $\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\subset L^{\infty}(\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X})))$ to cycles and filling chains in $\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\subset L^{\infty}(\mathcal{X})$.
The next step follows easily from the definition of the filling radius:
$\rho(\pi_{*}\omega;\mathcal{N}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X})))=\rho(\pi_{*}\omega;\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))).$
The final step is to apply Theorem 3.3 for the compact set $\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))$, to conclude that
$\rho(\pi_{*}\omega;\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X})))\leq\frac{1}{2}\operatorname{UW}_{k-1}(\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))).$
Assembling all these inequalities together, we obtain the result. ∎
[…]
Remark 3.16.
It follows that lifespans in dimensions $1$ and higher can be upper-bounded by $\operatorname{UW}_{0}(\operatorname{E}(\mathcal{X}))=\operatorname{diam}(\mathcal{X})$. It is easy to see that lifespans in dimension 00 are also bounded by $\operatorname{diam}(\mathcal{X})$.
Proof of Section 1.
Recall that
$d_{\omega}-b_{\omega}=2\lim_{r\to b_{\omega}+0}\rho(\omega_{r};\mathcal{N}_{r}(\mathcal{X}\subset L^{\infty}(\mathcal{X}))\subset L^{\infty}(\mathcal{X})),$
where $\omega_{r}$ maps to $\omega$ by the homology map induced by the inclusion. Upper-bounds for these filling radii are given by Theorem 3.15. To conclude, it remains to use the continuity property of the Urysohn width [Bal21, Theorem 2.4.1]:
$\lim_{r\to b_{\omega}+0}\operatorname{UW}_{k-1}(\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\operatorname{E}(\mathcal{X})))=\operatorname{UW}_{k-1}(\overline{\mathcal{N}}_{b_{\omega}}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))).$
3.4 Robust Bounds for Čech Lifespans via Treewidth
The estimates above have a drawback: There is no a priori way to relate the width of $\mathcal{X}\subset\mathbf{Y}$ and the width of its $r$-neighborhood $\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})$, while the bounds in Corollaries 3.121 and 1 depend on the width of a certain neighborhood of $\mathcal{X}$. By narrowing down the class of potential cores even further we can guarantee some bounds. For example, this happens if the core is an affine $k$-dimensional subspace (see Section 3.5). Here we introduce a modified version of treewidth that allows for bounds on lifespans in terms of $\mathcal{X}$ itself and not of its neighborhood. We start with an auxiliary definition that we will use to restrict the geometric complexity of potential cores.
Definition 3.17.
Let $(\mathbf{Y},\|\cdot\|)$ be a Banach space, and let $\mathcal{T}\subset\mathbf{Y}$ be a simplicial complex. Suppose there is a retraction map $f:\mathbf{Y}\to\mathcal{T}$. We say that $f$ is $C$-robust if for any finite set $\mathcal{X}\subset\mathbf{Y}$, and any $x\in\mathrm{conv}(\mathcal{X})$,
$\|f(x)-x\|\leq\max_{x^{\prime}\in\mathcal{X}}\|f(x^{\prime})-x^{\prime}\|+C\cdot\mathrm{rad}(\mathcal{X}\subset\mathbf{Y}).$
Remark 3.18.
Recall that a retraction $f:\mathbf{Y}\to\mathcal{T}$ is a continuous map fixing every point of $\mathcal{T}$. Since $\mathbf{Y}$ is contractible, the existence of a retraction implies that $\mathcal{T}$ is contractible.
Examples.
  1. 1.
    If $\mathcal{T}$ is a closed affine subspace of $\mathbf{Y}$, and $f$ is any linear projection on $\mathcal{T}$, then $f$ is 00-robust.
  2. 2.
    If $\mathcal{T}$ is nonempty closed convex subset of $\mathbf{Y}$, and $f$ is a continuous nearest-point projection (metric projection) on $\mathcal{T}$, then $f$ is 00-robust. [^10] It is known (see the discussion after [Meg12, Corollary 5.1.19]) that a nearest-point projection exists if $\mathbf{Y}$ is reflexive; but it does not have to be continuous in general.
  3. 3.
    If the assignment $x\mapsto\|f(x)-x\|$ is $C$-Lipschitz, then $f$ is $C$-robust.
  4. 4.
    If $\mathbf{Y}=\mathbb{R}^{2}$ is the Euclidean plane, and $\mathcal{T}$ is the union of $m\geq 3$ equiangular rays (or intervals) with a common endpoint, then there is a natural retraction to $\mathcal{T}$ which is $\frac{1}{\tan(\pi/m)}$-robust.
Definition 3.19 (Robust Treewidth).
Let $\mathcal{X}$ be a subset of a Banach space $(\mathbf{Y},\|\cdot\|)$. For an integer $k\geq 0$, and a real number $C\geq 0$, we say that the $C$-robust $k$-dimensional treewidth of $\mathcal{X}$ is at most $\delta$, if there is a $C$-robust retraction $f:\mathbf{Y}\to\mathcal{T}$ to a finite simplicial complex $\mathcal{T}\subset\mathbf{Y}$ of dimension at most $k$, displacing every point of $\mathcal{X}$ by distance at most $\delta$:
$\|f(x)-x\|\leq\delta\quad\forall x\in\mathcal{X}.$
The infimum of the numbers $\delta$ satisfying this condition will be denoted by $\operatorname{TW}^{C}_{k}(\mathcal{X}\subset\mathbf{Y})$.
It is immediate from the definition that $\operatorname{TW}^{C_{1}}_{k}(\mathcal{X}\subset\mathbf{Y})\geq\operatorname{TW}^{C_{2}}_{k}(\mathcal{X}\subset\mathbf{Y})$ for any $C_{1}\leq C_{2}$, and that $\operatorname{TW}^{C}_{k}(\mathcal{X}\subset\mathbf{Y})\geq\operatorname{TW}_{k}(\mathcal{X}\subset\mathbf{Y})$ for any $C$.
Theorem 3.20.
For any compact set $\mathcal{X}$ sitting in a Banach space $\mathbf{Y}$, any integer $k\geq 0$, any real numbers $C,r\geq 0$, and any homology class $\omega\in\operatorname{\tilde{H}}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y}))$,
$\rho(\omega;\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})\subset\mathbf{Y})\leq\operatorname{TW}^{C}_{k}(\mathcal{X}\subset\mathbf{Y})+Cr.$
Proof.
Suppose we are given a $C$-robust retraction $f:\mathbf{Y}\to\mathcal{T}$ to a $k$-dimensional simplicial complex $\mathcal{T}\subset\mathbf{Y}$, such that $\|x-f(x)\|\leq\delta$ for all $x\in\mathcal{X}$.
Let our homology class $\omega\in\operatorname{\tilde{H}}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y}))$ be represented by a singular cycle $S_{0}$ supported in $\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})$. Since $S_{0}$ is compact as a subset of $\mathbf{Y}$, it can be covered by finitely many open balls $B_{r}(x),x\in\mathcal{X}^{\prime}\subset\mathcal{X},|\mathcal{X}^{\prime}|<\infty$ (here $B_{r}(x)$ denotes the open ball of radius $r$ centered at $x$). Consider the union $U=\bigcup\limits_{x\in\mathcal{X}^{\prime}}B_{r}(x)$ and choose any partition of unity subordinate to the cover of $U$ by $B_{r}(x),x\in\mathcal{X}^{\prime}$. This partition of unity induces a map from $U$ to the nerve $Z$ of that cover. By the nerve theorem this map $\phi:U\to Z$ is a homotopy equivalence, with an evident homotopy inverse $\psi:Z\to U$, which sends every vertex of $Z$ to the center of the corresponding ball in the cover $U=\bigcup\limits_{x\in\mathcal{X}^{\prime}}B_{r}(x)$, and extends affinely on the rest of $Z$. Therefore, $\omega|_{U}=\psi_{*}\phi_{*}\omega|_{U}\in\operatorname{\tilde{H}}_{k}(U)$, and the class $\omega$ can be represented by a cycle $S$ supported in $\psi(Z)$; in other words, $S$ is obtained by gluing flat $k$-simplices with vertices in $\mathcal{X}$, and with the circumradius of each simplex at most $r$.
Now we continuously deform $S$ using the linear homotopy $h:S\times[0,1]\to\mathbf{Y}$ given by
$h(x,t)=(1-t)x+tf(x).$
By the robustness property of $f$, every point $x\in S$ is moved by a distance at most $\delta+Cr$, so the continuous deformation always stays in $\mathcal{N}_{\delta+Cr}(S\subset\mathbf{Y})$. At time $t=1$, the deformed cycle $h(S,1)$ lies in $\mathcal{T}$ and bounds a $(k+1)$ chain within $h(S,1)$, because $\mathcal{T}$ is of dimension $k$ with trivial $\operatorname{\tilde{H}}_{k}(\mathcal{T})$. Therefore, $S$ is nullhomologous in $\mathcal{N}_{\delta+Cr}(S\subset\mathbf{Y})$ (basically, $S$ bounds the $(k+1)$-dimensional trace of the homotopy $h$). ∎
[…]
Proof.
We again employ formula (2.3), now bounding filling radii by 3.20:
$d_{\omega}-b_{\omega}$ $=\lim_{r\to b_{\omega}+0}\rho(\omega_{r};\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})\subset\mathbf{Y})$
$\leq\lim_{r\to b_{\omega}+0}\operatorname{TW}^{C}_{k}(\mathcal{X}\subset\mathbf{Y})+Cr$
$=\operatorname{TW}^{C}_{k}(\mathcal{X}\subset\mathbf{Y})+Cb_{\omega}.$
For $b_{\omega}\geq 1$, if we divide the entire inequality by $b_{\omega}$, we get the result. ∎
Remark 3.21 (Multiplicative Persistence).
The corollary above implies a robust estimate on the multiplicative persistence (see [BKS17, AC22]) of intervals in the persistence barcode of the Čech filtration $\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y})$. Notice that on the right-hand side, there is no reference to the birth time $b_{\omega}$. Therefore, $C$-robust treewidth gives a strong bound for late-born ($b_{\omega}\geq 1$) homology classes in terms of multiplicative persistence. Note also that the proof above also gives the following estimates for $(b_{\omega},d_{\omega})$:
  • If $b_{\omega}=0$, then $d_{\omega}\leq\operatorname{TW}^{C}_{k}(\mathcal{X}\subset\mathbf{Y})$.
  • If $b_{\omega}\geq\operatorname{TW}^{C}_{k}(\mathcal{X}\subset\mathbf{Y})$, then $\dfrac{d_{\omega}}{b_{\omega}}\leq C+2$.
3.5 Robust Bounds for Čech Lifespans via Kolmogorov Width and ${}_{\infty}$
For a given compact set $\mathcal{X}$ in a Banach space $\mathbf{Y}$ (e.g., a point cloud in $\mathbb{R}^{N}$), the usual Principal Component Analysis (PCA) approximates $\mathcal{X}$ by affine $k$-planes through minimizing the $\ell^{2}$-distance from $\mathcal{X}$ to a plane. Here we consider an $\ell^{\infty}$-version of the classical procedure, by optimizing the $\ell^{\infty}$-distance instead. Recall that in the original PCA, the eigenvalues of the covariance matrix are called variances. This motivates the following definition.
Definition 3.22 (${}_{\infty}$).
[…]
$\nu_{k+1}(\mathcal{X}\subset\mathbf{Y})=\inf_{\mathcal{P}\in\mathcal{A}_{k}}\sup_{x\in\mathcal{X}}\mathbf{d}_{\mathbf{Y}}(x,\mathcal{P})$
where $\mathbf{d}_{\mathbf{Y}}(x,\mathcal{P})$ is the distance from $x$ to the $k$-subspace $\mathcal{P}$, i.e., $\mathbf{d}_{\mathbf{Y}}(x,\mathcal{P})=\inf_{p\in\mathcal{P}}\|x-p\|$.
Remark 3.23 (Kolmogorov Width).
[…]
Remark 3.24.
The $\infty$-variance is just a different name for Kolmogorov widths (cf. 3.6):
$\nu_{k+1}(\mathcal{X}\subset Y)=\operatorname{KW}_{k}(\mathcal{X}\subset\mathbf{Y}).$
Theorem 3.25.
Let $\mathcal{X}$ be a subset of a Banach space $\mathbf{Y}$. Then for any homology class $\omega\in\operatorname{\tilde{H}}_{k}(\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y}))$, $k\geq 0$,
$\rho(\omega;\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})\subset\mathbf{Y})\leq\nu_{k+1}(\mathcal{X}\subset\mathbf{Y})=\operatorname{KW}_{k}(\mathcal{X}\subset\mathbf{Y}).$
Proof.
Apply Theorem 3.20 with $\mathcal{T}$ being a $k$-dimensional affine subspace, and $f$ being a 00-robust nearest-point projection. ∎
[…]
Proof.
Combine formula (2.3) with 3.25. ∎
Remark 3.26.
Notice that the corollary above gives strong estimates for point clouds in $\mathbb{R}^{N}$. For example, if the variance $\nu_{k_{0}+1}(\mathcal{X})$ is small, then the corollary above implies that one can ignore homology classes in $\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$ for $k\geq k_{0}$. We should note that, by its very definition, ${}_{\infty}$ is not as robust as the original PCA against outliers. We discuss this difference and pose a statistical question about the relationship between original PCA and persistent homology in Section 6.
Remark 3.27.
Notice that all lifespan bounds in Corollaries 3.12, 1, 1, 1, and 1 are stated in terms of homological lifespans ($d_{\omega}-b_{\omega}$). 2.22 implies that all these bounds therefore apply to the length of every interval in either $\mathrm{PD}_{k}(\mathcal{X})$ or $\widecheck{\mathrm{PD}}_{k}(\mathcal{X}\subset\mathbf{Y}).$
4 Bounding Lifespans by Spread
This section is concerned with global upper bounds for Čech lifespans of homology classes in arbitrary degrees dimension. Our results in this section are somewhat parallel to the following result for VR lifespans.
Theorem 4.1 (VR-lifespans via Spread; [LMO23, Proposition 9.3 and Remark 9.8]).
Let $\mathcal{X}$ be a compact metric space and let $\omega\in\mathrm{Spec}_{k}(\mathcal{V}_{\bullet}(\mathcal{X}))$. Then $d_{\omega}-b_{\omega}\leq\operatorname{spread}(\mathcal{X})$.
Here Katz’s notion of spread is used [Kat83]. By definition, $\operatorname{spread}(\mathcal{X})$ is the infimum of the real numbers $\delta\geq 0$ for which there is a finite subset $A=\{a_{1},\ldots,a_{m}\}\subset\mathcal{X}$ with $\operatorname{diam}(A)\leq\delta$ and $\mathbf{d}^{\mathcal{X}}_{\mathrm{H}}(\mathcal{X},A)\leq\delta$.
In the rest of this section, we work in a Banach space $(\mathbf{Y},\|\cdot\|)$ and, for each compact set $\mathcal{X}\subset\mathbf{Y}$, we discuss a certain variation of the spread for which we prove an analogue of 4.1.
Definition 4.2 (Übercontractibility).
A set $\mathcal{T}$, sitting in a Banach space $(\mathbf{Y},\|\cdot\|)$, will be called $\delta$-übercontractible if its neighborhood $\mathcal{N}_{r}(\mathcal{T}\subset\mathbf{Y})$ is contractible for every $r\geq\delta$.
Remark 4.3 (Acyclicity).
For our purposes we actually need a slightly weaker condition: all neighborhoods $\mathcal{N}_{r}(\mathcal{T})$, $r\geq\delta$, are acyclic (have trivial homology groups) rather than contractible (have trivial homotopy groups). The acyclicity condition is equivalent to having almost trivial Čech persistence diagram (trivial beyond death time $\delta$), and is implied by übercontractibility as defined above.
For constructive examples of übercontractible sets, we refer the reader to Appendix B, where we show that the cut-locus of a convex set is typically 00-übercontractible.
Definition 4.4 (Überspread).
For a compact set $\mathcal{X}\subset\mathbf{Y}$, we define its überspread as
[…]
Figure 3: The bound $d-b\leq 2\,\ddot{u}$ in Section 1 is sharp; here $\mathbf{Y}=\mathbb{R}^{2}$ and $\ddot{u}=\operatorname{{\ddot{u}}-spread}(\mathcal{X}\subset\mathbb{R}^{2})$.
Proof.
Via the functorial nerve theorem (2.6), we will argue at the level of neighborhood filtrations, as opposed to simplicial filtrations. Let $\mathcal{T}$ be a $\delta$-übercontractible simplicial complex in $\mathbf{Y}$ which is Hausdorff distance at most $\delta$ away from $\mathcal{X}$, where $\delta$ is just a tiny bit greater than $\operatorname{{\ddot{u}}-spread}(\mathcal{X}\subset\mathbf{Y})$. Then, for all $r>0$ we have
[…]
where $i_{r,r+2\delta}$, $p_{r}$, and $q_{r+\delta}$ are the obvious inclusion maps. Applying the homology functor to the above diagram gives, for all $r>0$, that
$\operatorname{\tilde{H}}_{k}(i_{r,r+2\delta})=\operatorname{\tilde{H}}_{k}(p_{r})\circ\operatorname{\tilde{H}}_{k}(q_{r+\delta})=0$
since $\mathcal{N}_{r+\delta}(\mathcal{T}\subset\mathbf{Y})$ is contractible. The this means that $d_{\omega}-b_{\omega}\leq 2\delta$. The proof follows. ∎
Remark 4.5 (Katz spread vs. $\operatorname{{\ddot{u}}-spread}$).
Let us compare Section 1 with [LMO23, Proposition 9.3]. Suppose we are given a compact metric space $\mathcal{X}$ and a finite subset $A=\{a_{1},\ldots,a_{m}\}\subset\mathcal{X}$ with $\operatorname{diam}(A)\leq\delta$ and $\mathbf{d}^{\mathcal{X}}_{\mathrm{H}}(\mathcal{X},A)\leq\delta$. In other words, the Katz spread of $\mathcal{X}$ does not exceed $\delta$:
$\operatorname{spread}(\mathcal{X})\leq\delta.$
Embed $\mathcal{X}$ in $L^{\infty}(\mathcal{X})$ via the Kuratowski map $\kappa$, and note that the set $\kappa(A)\subset L^{\infty}(\mathcal{X})$ is $\delta/2$-übercontractible. Hence, we can apply Section 1 and conclude that Čech lifespans of $\kappa(\mathcal{X})\subset L^{\infty}(\mathcal{X})$ do not exceed $2\delta$. Therefore, VR lifespans of $\mathcal{X}$ do not exceed $4\delta$. This is weaker than the conclusion of [LMO23, Proposition 9.3], which tells us that in this situation VR lifespans of $\mathcal{X}$ do not exceed $\operatorname{spread}(\mathcal{X})\leq\delta$.
In general, the estimate in Section 1 cannot be improved (see Figure 3). However, the factor of $2$ in that estimate can sometimes be removed if the comparison set $\mathcal{T}$ admits a $1$-Lipschitz nearest-point projection from $\mathbf{Y}$.
Theorem 4.6.
Let $\mathcal{X}$ be a compact subset of a Banach space $(\mathbf{Y},\|\cdot\|)$. Assume that $\mathrm{conv}(\mathcal{X})$ admits a $1$-Lipschitz nearest-point projection to a closed convex set $\mathcal{T}\subset\mathbf{Y}$ [^11] For example, this is true whenever $\mathbf{Y}$ is Euclidean. Then any $\omega\in\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$, $k\geq 0$, has lifespan at most $\mathbf{d}_{\mathrm{H}}^{\mathbf{Y}}(\mathcal{X},\mathcal{T})$.
Proof.
Denote $\delta=\mathbf{d}_{\mathrm{H}}^{\mathbf{Y}}(\mathcal{X},\mathcal{T})$. By assumption, there exists a $1$-Lipschitz nearest-point projection $\pi:\mathrm{conv}(\mathcal{X})\rightarrow\mathcal{T}$; there is also a (possibly discontinuous) map $j:\mathcal{T}\to\mathcal{X}$ such that
$\|x-\pi(x)\|\leq\delta\,\,\mbox{and}\,\,\|a-j(a)\|\leq\delta\,\,\mbox{$\forall x\in\mathcal{X}$ and $a\in\mathcal{T}$}.$
It is easy to see that $\pi$ and $j$ induce, for each $\epsilon>0$, simplicial maps
$\pi_{\epsilon}:\mathcal{C}_{\epsilon}(\mathcal{X}\subset\mathbf{Y})\rightarrow\mathcal{C}_{\epsilon}(\mathcal{T}\subset\mathbf{Y})\,\,\mbox{and}\,\,j_{\epsilon}:\mathcal{C}_{\epsilon}(\mathcal{T}\subset\mathbf{Y})\rightarrow\mathcal{C}_{\epsilon+\delta}(\mathcal{X}\subset\mathbf{Y}).$
To see this in the case of $\pi$ we proceed as follows. Let $\sigma=\{x_{0},\ldots,x_{n}\}\in\mathcal{C}_{\epsilon}(\mathcal{X}\subset\mathbf{Y}).$ Let $y_{\sigma}\in\mathrm{conv}(\mathcal{X})$ satisfy $\|x_{i}-y_{\sigma}\|\leq\epsilon$ for all $i$. Then, since $\pi:\mathrm{conv}(\mathcal{X})\to\mathcal{T}$ is 1-Lipschitz, we have
$\|\pi(x_{i})-\pi(y_{\sigma})\|\leq\|x_{i}-y_{\sigma}\|\leq\epsilon$
for all $i$. Hence, $\pi_{\epsilon}(\sigma)\in\mathcal{C}_{\epsilon}(\mathcal{T}\subset\mathbf{Y}).$
In the case of $j$, let $\sigma=\{x_{0},\ldots,x_{n}\}\in\mathcal{C}_{\epsilon}(\mathcal{T}\subset\mathbf{Y}).$ Let $y_{\sigma}\in\mathbf{Y}$ satisfy $\|x_{i}-y_{\sigma}\|\leq\epsilon$ for all $i$. Then,
$\|j(x_{i})-y_{\sigma}\|\leq\|j(x_{i})-x_{i}\|+\|x_{i}-y_{\sigma}\|\leq\delta+\epsilon$
for all $i$. Hence, $j_{\epsilon}(\sigma)\in\mathcal{C}_{\epsilon+\delta}(\mathcal{X}\subset\mathbf{Y}).$
We then have the following (not-necessarily commutative) diagram:
[…]
where $i_{\epsilon,\epsilon+\delta}$ is the obvious inclusion map.
Claim: For each $\epsilon>0$ the maps $j_{\epsilon}\circ\pi_{\epsilon}$ and $i_{\epsilon,\epsilon+\delta}$ are contiguous.
To see the claim, let $\sigma=\{x_{0},\ldots,x_{n}\}$ be any simplex in $\mathcal{C}_{\epsilon}(\mathcal{X}\subset\mathbf{Y})$. We will prove that $\tau:=\sigma\cup j\circ\pi(\sigma)$ is a simplex in $\mathcal{C}_{\epsilon+\delta}(\mathcal{X}\subset\mathbf{Y}).$ This requires us to find a point $y_{\tau}\in\mathbf{Y}$ such that $\|v-y_{\tau}\|\leq\epsilon+\delta$ for all $v\in\tau$.
Let $y_{\sigma}\in\mathbb{R}^{d}$ be such that $\|x_{i}-y_{\sigma}\|\leq\epsilon$ for all $i$. Notice that then we can write
$\tau=\{x_{0},\ldots,x_{n}\}\cup\{j\circ\pi(x_{0}),\ldots,j\circ\pi(x_{n})\}.$
Then, we let $y_{\tau}:=\pi(y_{\sigma})$ and calculate
$\max_{v\in\tau}\|v-y_{\tau}\|=\max\left(\max_{i}\|x_{i}-\pi(y_{\sigma})\|,\max_{i}\|j\circ\pi(x_{i})-\pi(y_{\sigma})\|\right).$
For the first argument of the maximum above,
$\|x_{i}-\pi(y_{\sigma})\|$ $\leq\|x_{i}-\pi(x_{i})\|+\|\pi(x_{i})-\pi(y_{\sigma})\|$
$\leq\|x_{i}-\pi(x_{i})\|+\|x_{i}-y_{\sigma}\|$
$\leq\delta+\epsilon.$
For the second argument,
$\|j\circ\pi(x_{i})-\pi(y_{\sigma})\|$ $\leq\|j\circ\pi(x_{i})-\pi(x_{i})\|+\|\pi(x_{i})-\pi(y_{\sigma})\|$
$\leq\|j\circ\pi(x_{i})-\pi(x_{i})\|+\|x_{i}-x_{\sigma}\|$
$\leq\delta+\epsilon,$
which establishes the claim.
Now, going back to the proof of the theorem: apply the homology functor $\operatorname{\tilde{H}}_{k}$ to the diagram above and obtain the following commutative diagram:
[…]
Now, $\operatorname{\tilde{H}}_{k}(\mathcal{C}_{\epsilon}(\mathcal{T}\subset\mathbf{Y}))=0$ since $\mathcal{T}$ is convex. Then,
$\operatorname{\tilde{H}}_{k}(i_{\epsilon,\epsilon+\delta})=\operatorname{\tilde{H}}_{k}(j_{\epsilon})\circ\operatorname{\tilde{H}}_{k}(\pi_{\epsilon})=0,$
which completes the proof. ∎
5 Bounding Extinction Times
In Section 3, we discussed bounds for the lifespans of individual homology classes in degree $\geq k$ via different notions of $k$-width from metric geometry. However, as Figure 2 suggests, these $k$-widths do not give any bound for the lower homology classes in lower degrees. In other words, if $\operatorname{AW}_{k}(\mathcal{M}\subset\mathbb{R}^{N})$ (or another $k$-width) is small, the filling radius (or lifespan) of a homology class $\alpha\in\operatorname{\tilde{H}}_{j}(\mathcal{M})$ for $j<k$ can still be very large. It is easy to see that the example in Figure 2 can be generalized to any dimension and codimension. In particular, $k$-widths do not say much about the size of topological features in dimensions lower than $k$.
In this section, we use a different approach and propose a global bound to the lifespans of homology classes in all degrees at once. Our goal is to bound the following quantity:
Definition 5.1 (Extinction Time).
Let $\mathcal{X}$ be a compact metric space. Then,
$\xi(\mathcal{X})=\sup_{k\geq 0}\sup\{d_{\omega}\mid\omega\in\mathrm{Spec}_{k}(\mathcal{V}_{\bullet}(\mathcal{X}))\}$
is called the VR-extinction time of $\mathcal{X}$. Similarly, define $\widecheck{\xi}(\mathcal{X}\subset\mathbf{Y})$ as the Čech-extinction time for Čech filtration in a Banach space $\mathbf{Y}$. We will use the terms VR-extinction (respectively Čech-extinction) for short.
Notice that with the VR-extinction $\xi(\mathcal{X})$ time, we are not only bounding the lifespans of homological features with birth time $=0$, but we are bounding the death time of all homological features (in degree $1$ and higher) appearing throughout the filtration. Also, note that the radius of $\mathcal{X}$ (2.35) automatically gives an upper bound for VR-extinction time, i.e. $\xi(\mathcal{X})\leq\mathrm{rad}(\mathcal{X})$ by [LMO23, Remark 9.4]). A similar statement is clearly true for Čech-extinction time. i.e., $\widecheck{\xi}(\mathcal{X}\subset\mathbf{Y})\leq\mathrm{rad}(\mathcal{X}\subset\mathbf{Y})$.
Remark 5.2.
Due to 2.21, any bound on extinction times of a filtration will result in an upper bound on the right endpoints of any interval in the barcode of said filtration; cf. 2.22.
In the following, we aim to give much finer estimates for extinction times in both the VR and Čech settings.
Remark 5.3 (Motivating Example).
Here, we give a toy example to motivate the notion of extinction defined above. Let $\mathrm{E}$ be the $(N-1)$-dimensional ellipsoid in $\mathbb{R}^{N}$ given by
$\mathrm{E}=\left\{\mathbf{x}\in\mathbb{R}^{N}~{}\middle|~{}\sum_{i=1}^{N}\dfrac{x_{i}^{2}}{a_{i}^{2}}=1\right\},$
where $a_{1}>a_{2}>\dots>a_{N}>0$. While $\mathrm{E}$ has trivial homology groups in low degrees, one can easily add some topology to $\mathrm{E}$ by adding $k$-handles for $10<k<N-$ 1 as follows. Fix $10<k<N-$ 1. Let
$\overline{\mathrm{E}}=\left\{\mathbf{x}\in\mathbb{R}^{N}~{}\middle|~{}\sum_{i=1}^{N}\dfrac{x_{i}^{2}}{a_{i}^{2}}\leq 1\right\},$
be the solid ellipsoid. Given an integer $m\geq 1$, we will use to notation $\overline{\mathrm{E}}^{m}=\{\mathbf{x}\in\overline{\mathrm{E}}\mid x_{i}=0\mbox{ for }i>m\}$. Let $\{D_{1},D_{2},\dots,D_{\ell_{k}}\}$ be $\ell_{k}$ disjoint small disks in $\overline{\mathrm{E}}^{N-k}$. Then, consider the following surgery operation. For $1\leq j\leq\ell_{k}$, let $\Omega_{j}=\overline{\mathrm{E}}\cap(D_{j}\times\mathbb{R}^{k})$. Let $S_{j}=\mathrm{E}\cap\Omega_{j}$, and $T_{j}=\operatorname{cl}(\partial\Omega_{j}\setminus S_{j})$. Then, swapping the $S_{j}$ and the $T_{j}$ will give a new closed manifold
$\widehat{\mathrm{E}}=\left(\mathrm{E}\setminus\bigcup_{j=1}^{\ell_{k}}S_{j}\right)\cup\bigcup_{j=1}^{\ell_{k}}T_{j}\subset\mathbb{R}^{N}.$
See Figure 4 for an illustration. While $\operatorname{\tilde{H}}_{k}(\mathrm{E})$ is trivial for $10<k<N-$ 1, $\mbox{rank}(\operatorname{\tilde{H}}_{k}(\widehat{\mathrm{E}}))\geq{\ell_{k}}$ because of the homology classes generated by the $k$-handles $\{T_{j}\}$. By choosing $\{D_{j}\}$ such that $\{\Omega_{j}\}$ are all pairwise disjoint, one can obtain $\widehat{\mathrm{E}}$ with nontrivial homology in all desired degrees $k$.
Figure 4: The space $\widehat{\mathrm{E}}$ from 5.3 for the case $N=3$ and $k=1$.
Notice that, through Kolmogorov widths, Section 1 yields that
$d_{\omega}-b_{\omega}\leq\operatorname{KW}_{k}(\widehat{\mathrm{E}}\subset\mathbb{R}^{N})=a_{k+1}$
for any class $\omega\in\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathrm{R}^{N})).$ By suitably choosing the numbers $\{a_{i}\}$, all $k$-widths for $k\leq\dim\widehat{\mathrm{E}}-1$ can be made arbitrarily large. However, it is not hard to see that, in this particular example, the death times of all these homology classes do not exceed the smallest axis length $a_{N}$. This example therefore shows that $k$-widths can highly overestimate lifespans. In other words, we have that the Čech extinction satisfies $\widecheck{\xi}(\widehat{\mathrm{E}}\subset\mathbb{R}^{N})\leq a_{N}$. In the following sections, we estimate extinction radii by comparing a given space to a nearby topologically trivial space.
5.1 Bounding Čech-extinction with Convex Hulls
Definition 5.4 (Convexity Deficiency).
For a compact set $\mathcal{X}$ sitting in a Banach space $(\mathbf{Y},\|\cdot\|)$, we define the convexity deficiency of $\mathcal{X}$ as
$\operatorname{cdef}(\mathcal{X}\subset\mathbf{Y})=\mathbf{d}_{\mathrm{H}}^{\mathbf{Y}}(\mathcal{X},\mathrm{conv}(\mathcal{X}))=\sup_{y\in\mathrm{conv}(\mathcal{X})}\inf_{x\in\mathcal{X}}\|y-x\|.$
[…]
Proof.
For any $r\geq 0$, and any points $x_{1},\ldots,x_{m}\in\mathcal{X}$, the intersection of balls $\bigcap_{i=1}^{m}B_{r}(x_{i})$ is nonempty in $\mathbf{Y}$ if and only if the intersection $\bigcap_{i=1}^{m}(B_{r}(x_{i})\cap\mathrm{conv}(\mathcal{X}))$ is non-empty. Therefore, the filtered Čech complexes $\mathcal{C}(\mathcal{X}\subset\mathbf{Y})$ and $\mathcal{C}(\mathcal{X}\subset\mathrm{conv}(\mathcal{X}))$ are identical. However, the homotopy type of $\mathcal{C}_{r}(\mathcal{X}\subset\mathrm{conv}(\mathcal{X}))$ is trivial whenever $r\geq\operatorname{cdef}(\mathcal{X}\subset\mathbf{Y})$, so no topological feature persists beyond time $\operatorname{cdef}(\mathcal{X}\subset\mathbf{Y})$. ∎
Remark 5.5 (Stability vs. Extinction).
The stability theorems (Section 2.1.3) do not imply the previous result as they would only yield bounds on lifespans but not on death times. Indeed, 2.16 implies that
$\mathbf{d}_{b}(\widecheck{\mathrm{PD}}_{k}(\mathcal{X}\subset\mathbf{Y}),\widecheck{\mathrm{PD}}_{k}(\mathrm{conv}(\mathcal{X})\subset\mathbf{Y})))\leq\mathbf{d}_{\mathrm{H}}^{\mathbf{Y}}(\mathcal{X},\mathrm{conv}(\mathcal{X}))=\operatorname{cdef}(\mathcal{X}\subset\mathbf{Y}).$
Since $\widecheck{\mathrm{PD}}_{k}(\mathrm{conv}(\mathcal{X})\subset\mathbf{Y})=\emptyset$, the stability result and the definition of the bottleneck distance [EH10] imply that for any $(b,d)\in\widecheck{\mathrm{PD}}_{k}(\mathcal{X}\subset\mathbf{Y})$,
$d-b\leq 2\operatorname{cdef}(\mathcal{X}\subset\mathbf{Y}).$
However, both $b$ and $d$ can be arbitrarily large. Meanwhile, Section 1 bounds elements of $\widecheck{\mathrm{PD}}_{k}(\mathcal{X}\subset\mathbf{Y})$ independently of $k$: for any $(b,d)\in\widecheck{\mathrm{PD}}_{k}(\mathcal{X}\subset\mathbf{Y})$,
$d\leq\operatorname{cdef}(\mathcal{X}\subset\mathbf{Y}).$
Furthermore:
  • Section 1 applies to all elements of $\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$ – not just those that can be associated to an interval in $\widecheck{\mathrm{PD}}_{k}(\mathcal{X}\subset\mathbf{Y})$; see 5.2.
  • The stability theorems do not in general apply in order to obtain bounds for either the lifetime or extinction of arbitrary classes in $\mathrm{Spec}_{k}(\mathcal{V}_{\bullet}(\mathcal{X}))$ or $\mathrm{Spec}_{k}(\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbf{Y}))$. For example, the Filling Radius of a manifold is not stable under the Gromov-Hausdorff distance; see [LMO23, Sections 9.4 and 9.5].
Remark 5.6.
Example in Figure 3 shows that it is not true that the extinction $\widecheck{\xi}(\mathcal{X}\subset\mathbf{Y})$ is upper-bounded by the überspread $\operatorname{{\ddot{u}}-spread}(\mathcal{X}\subset\mathbf{Y})$.
While the above remark shows that the distance to the nearest übercontractible space (überspread) fails to bound extinction, it might be still true that distance to a specific cleverly chosen übercontractible space might give an estimate on extinction. One natural choice for the role of an übercontractible space $\mathcal{T}$ approximating $\mathcal{X}\subset\mathbf{Y}$ is the convex hull of $\mathcal{X}$; this is exactly what has just been discussed; with this choice, there is an extinction bound (Theorem 1), but it is tempting to improve it by choosing a finer $\mathcal{T}$. Another natural choice of $\mathcal{T}$ is given by the cut-locus of the boundary of the convex hull of $\mathcal{X}$ (see the discussion in Appendix B). Unfortunately, the extinction time is not bounded from above by the distance from $\mathcal{X}$ to this $\mathcal{T}$ (chosen as the cut-locus of $\partial\mathrm{conv}(\mathcal{X})$, assuming $\mathrm{conv}(\mathcal{X})$ full-dimensional). We omit the discussion of examples in view of the negative nature of the result, but they can be obtained as subsets of a square (one can take a square and cut out a large off-centered disk).
Question 5.7 (Čech extinction and cut-locus of $\mathrm{conv}(\mathcal{X})$).
Let $\mathcal{X}$ be a compact subset of a Banach space $(\mathbf{Y},\|\cdot\|)$. Let $\mathcal{C}^{\partial\;\mathrm{conv}(\mathcal{X})}$ be the cut-locus of the convex hull of $\mathcal{X}$ (see Appendix B). Is it possible to upper-bound the extinction $\widecheck{\xi}(\mathcal{X}\subset\mathbf{Y})$ in terms of the Hausdorff distance to $\mathcal{C}^{\partial\;\mathrm{conv}(\mathcal{X})}$? i.e. for some $C_{\mathbf{Y}}>0$, is it true that
$\widecheck{\xi}(\mathcal{X}\subset\mathbf{Y})\leq C_{\mathbf{Y}}\cdot\mathbf{d}_{\mathrm{H}}^{\mathbf{Y}}(\mathcal{X},\mathcal{C}^{\partial\;\mathrm{conv}(\mathcal{X})})$
5.2 Bounding VR-extinction via the Tight Span
To get a bound on VR-extinctions, Theorem 1 can be applied for the Kuratowski distance-preserving embedding of $\mathcal{X}$ in $L^{\infty}(\mathcal{X})$, but the result will be sub-optimal. We refine it now.
The following definition is analogous to Definition 5.4 in that the convex hull of $\mathcal{X}\subset\mathbf{Y}$ is supplanted by $\operatorname{E}(\mathcal{X})$ (see 2.27). Recall that $\mathcal{X}$ naturally embeds in its tight span, so we can assume that $\mathcal{X}\subset\operatorname{E}(\mathcal{X})$.
Definition 5.8 (Hyperconvexity Deficiency).
The hyperconvexity deficiency of a compact metric space $\mathcal{X}$ is defined as the number
$\operatorname{hcdef}(\mathcal{X}):=\mathbf{d}_{\mathrm{H}}^{\operatorname{E}(\mathcal{X})}(\mathcal{X},\operatorname{E}(\mathcal{X})).$
The following corollary to 2.7 is analogous to Theorem 1.
Corollary 5.9 (Bounding VR Extinction).
Let $\mathcal{X}$ be a compact metric space. Then,
$\xi(\mathcal{X})\leq 2\,\operatorname{hcdef}(\mathcal{X}).$
Furthermore, this bound is tight (see Remark 5.10 below).
Proof.
The claim follows from 2.7 and 2.8 together with the facts that (1) $\mathcal{N}_{t}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))=\operatorname{E}(\mathcal{X})$ for all $t\geq\operatorname{hcdef}(\mathcal{X})$ and (2) $\operatorname{E}(\mathcal{X})$ is contractible. ∎
Remark 5.10 (Tightness of the bound).
Let $\mathcal{X}$ be the unit $\ell^{\infty}$ sphere in $\mathbb{R}^{2}$. Then, in that case, by results of [KK16], $\operatorname{E}(\mathcal{X})\cong\Big{(}[-1,1]\times[-1,1],\ell^{\infty}\Big{)}$ and we compute that $\operatorname{hcdef}(\mathcal{X})=1$. Since $\mathcal{N}_{t}(\mathcal{X}\subset\operatorname{E}(\mathcal{X}))\simeq\mathcal{X}\simeq S^{1}$ for every $10<t<$ 1, by Theorem 2.7 we see that, for $\omega=[S^{1}]$, $d_{\omega}=2$ (see also [LMO23, Corollary 7.4].
Recall from [LMO23, Proposition 9.1] that $\mathcal{V}_{r}(\mathcal{X})$ becomes contractible as soon as $r\geq\mathrm{rad}(\mathcal{X})$. The next proposition proves that the bound in Corollary 5.9 is never worse than this bound.
Proposition 5.11.
The inequality
$\operatorname{hcdef}(\mathcal{X})\leq\mathrm{rad}(\mathcal{X})$
hold for every compact metric space $\mathcal{X}$.
Remark 5.12 (Comparison of $\mathrm{rad}(\mathcal{X})$ and $\operatorname{hcdef}(\mathcal{X})$).
The upper bound given by Corollary 5.9 can be much smaller than $\mathrm{rad}(\mathcal{X})$. Indeed, Let $\mathcal{X}$ be any metric tree, then, in that case (by item (3) of Proposition 2.28) $\operatorname{E}(\mathcal{X})=\mathcal{X}$ so that $\operatorname{hcdef}(\mathcal{X})=0$. However, $\mathcal{X}$ can be chosen so that $\mathrm{rad}(\mathcal{X})$ (and also its spread) are arbitrarily large.
Proof of Proposition 5.11.
Assume that $\delta>\mathrm{rad}(\mathcal{X})$ and let $x_{0}\in\mathcal{X}$ be a point such that $\mathbf{d}_{\mathcal{X}}(x_{0},x)<\delta$ for all $x\in\mathcal{X}$. Recall that, according to equation (4), we have that
$f(x_{0})=\max_{x^{\prime}\in\mathcal{X}}\big{(}\mathbf{d}_{\mathcal{X}}(x_{0},x^{\prime})-f(x^{\prime})\big{)}.$
Let $x^{\prime}_{0}\in\mathcal{X}$ be such that $f(x_{0})=\mathbf{d}_{\mathcal{X}}(x_{0},x^{\prime}_{0})-f(x_{0}^{\prime})$. Notice that then, for a given for $f\in\operatorname{E}(\mathcal{X})$, by equation (4) we have both
$\|f-\mathbf{d}_{\mathcal{X}}(x_{0},\cdot)\|_{\infty}=f(x_{0})\,\,\mbox{and}\,\,\|f-\mathbf{d}_{\mathcal{X}}(x_{0}^{\prime},\cdot)\|_{\infty}=f(x_{0}^{\prime}).$
Adding these two expressions we obtain that
$\|f-\mathbf{d}_{\mathcal{X}}(x_{0},\cdot)\|_{\infty}+\|f-\mathbf{d}_{\mathcal{X}}(x_{0}^{\prime},\cdot)\|_{\infty}=f(x_{0})+f(x_{0}^{\prime})=\mathbf{d}_{\mathcal{X}}(x_{0},x_{0}^{\prime})<\delta.$
From this, we conclude that
$\min\big{(}\|f-\mathbf{d}_{\mathcal{X}}(x_{0},\cdot)\|_{\infty},\|f-\mathbf{d}_{\mathcal{X}}(x_{0}^{\prime},\cdot)\|_{\infty}\big{)}\leq\frac{\delta}{2}.$
Since $f\in\operatorname{E}(\mathcal{X})$ was arbitrary, this establishes that $\operatorname{hcdef}(\mathcal{X})\leq\frac{\delta}{2}$ from which the claim follows. ∎
6 Final Remarks
Here we provide some remarks that could suggest further exploration.
Widths and Lifespans.
While our results providing bounds on lifespans via widths are of theoretical nature, they can be quite useful in applications in the following way. Computing the exact Alexandrov, Urysohn, or Kolmogorov $k$-width for a given set $\mathcal{X}$ can be computationally challenging, but note that by definition, these widths are basically infima over some measurements related to a $k$-dimensional space. While finding this optimal $k$-dimensional space is very hard, any relevant measure taken over a meaningful and well-chosen $k$-dimensional space gives an upper bound for these lifespans. Therefore, while computing the optimal such bound may not be feasible, these results can be effectively used to give some rough upper bounds for the lifespans of relevant topological features.
We would like to add that while we bring ideas from metric geometry to applied algebraic topology, our results can be used in the reverse direction, too. In particular, the lifespans are upper bounded by various width notions in Section 3, which implies that the lifespans gives lower bounds for the width estimation. In other words, for a given metric space $\mathcal{X}$, the maximum lifespan of a class in $\mathrm{Spec}_{k}(\mathcal{V}_{\bullet}(\mathcal{X}))$ gives a lower bound for the relevant $k$-width (e.g., $\operatorname{AW}_{k}(\mathcal{X}),\operatorname{UW}_{k}(\mathcal{X})$), which can be used as an effective computational approach to estimate $k$-widths.
Original PCA vs. Lifespans.
In Section 3.5, we give bounds for Čech lifespans via the variances $\{\nu_{k}\}$ induced by ${}_{\infty}$ (Section 1) As one can easily notice, in order to have such a rigorous bound, we modified the usual PCA definition and considered an $\ell^{\infty}$-variant. However, this makes the ${}_{\infty}$ structure highly sensitive to outliers, in contrast with the original PCA (with $\ell^{2}$-norm). On the other hand, standard PCA is a mainstream, highly effective dimension reduction tool for real-life applications with several computational techniques available. While our results do not say anything about the relation between the original PCA and PH, an experimental result relating the lifespans of bars in the persistence diagram of the Čech filtration $\mathcal{C}_{\bullet}(\mathcal{X}\subset\mathbb{R}^{N})$ with variances $\widetilde{\nu}_{k+1}(\mathcal{X})$ of original PCA would be very interesting, for a given finite set of points $\mathcal{X}\subset\mathbb{R}^{N}$. It would particularly interesting and useful for real like applications to to carry out a statistical comparative analysis for random finite subsets $\mathcal{X}$ in $\mathbb{R}^{N}$ (and for $k\geq 1$).
Principal Curves and Surfaces.
One can notice that when describing the width based arguments in Section 3, we first introduce a $k$-dimensional optimal core $\Gamma_{k}$ for a given set $\mathcal{X}$, then the $k$-width $\mathrm{W}_{k}(\mathcal{X})$ is defined as some kind of “distance” from $\mathcal{X}$ to $k$-core $\Gamma_{k}$. Hence, when $\mathrm{W}_{k}(\mathcal{X})$ is small, in metric geometry, $\mathcal{X}$ is considered as “essentially $k$-dimensional”. There is a similar notion in statistics called Principal Curves and Surfaces [HS89, Del01, OE11]. While principal curves and surfaces are defined as $1$- and $2$-dimensional objects, one can easily generalize the idea to any dimension $k$, e.g. principal $k$-manifolds. In our setting, for a compact subset $\mathcal{X}$ in a Banach space $\mathbf{Y}$, principal curves and surfaces can be considered as $k$-dimensional objects $\Sigma_{k}$ which minimizes $\ell^{2}$-distance from $\mathcal{X}$ to $\Sigma_{k}$ for $k=1,2$ with some normalization condition on $\Sigma_{k}$. In this respect, one can see principal curves and surfaces as a nonlinear generalization of principal component analysis (PCA).
In that respect, for a given set $\mathcal{X}$, our $k$-cores and principal $k$-manifolds can be considered as analogous objects. In Section 3.5, we defined ${}_{\infty}$ as $\ell^{\infty}$-version of the original $\ell^{2}$ PCA . Similarly, our $k$-cores $\Gamma_{k}$ minimize the $\ell^{\infty}$-distance between $\mathcal{X}$ and $\Gamma_{k}$ while principal $k$-manifolds $\Sigma_{k}$ minimizes the $\ell^{2}$-distance between $\mathcal{X}$ and $\Sigma_{k}$. By using this analogy, as principal curves and surfaces are suggested as dimension reduction method, one can consider our $k$-cores represent the essential structure of $\mathcal{X}$ when $\mathrm{W}_{k}(\mathcal{X})$ is small. Furthermore, just like the previous part (PCA vs ${}_{\infty}$), it would be interesting to make a statistical study of the relation between the Čech lifespans and $\ell^{2}$-distance to principal $k$-manifolds.
Appendix A Auxiliary Properties of Widths
It was mentioned in Section 3.1 that
$\operatorname{AW}_{k}(\mathcal{X}\subset\mathbf{Y})\leq\operatorname{UW}_{k}(\mathcal{X})\leq 2\operatorname{AW}_{k}(\mathcal{X}\subset\mathbf{Y}).$
The right-hand side inequality is trivial, whereas the inequality on the left requires an explanation.
Lemma A.1 (Alexandrov [Ale33]).
For any compact set $\mathcal{X}$ in a Banach space $\mathbf{Y}$,
$\operatorname{AW}_{k}(\mathcal{X}\subset\mathbf{Y})\leq\operatorname{UW}_{k}(\mathcal{X}).$
Proof.
Suppose $\operatorname{UW}_{k}(\mathcal{X})<\delta$, and let us show that $\operatorname{AW}_{k}(\mathcal{X}\subset\mathbf{Y})<\delta$. There is a continuous map $f:\mathcal{X}\to\Delta^{k}$ to a $k$-dimensional complex with fibers of diameter $<\delta$. Subdivide $\Delta^{k}$ very finely, so that the preimage of any open star of $\Delta^{k}$ under the map $f$ has diameter $<\delta$ [^12] Recall that the open star $S_{v}$ of a vertex $v\in\Delta^{k}$ is the union of the relative interiors of all simplices of $\Delta^{k}$ that contain $v$. For each vertex $v\in\Delta^{k}$, pick a point $c_{v}$ in the preimage of the open star $S_{v}$ (unless this preimage is empty; in this case, we can safely remove $S_{v}$ from $\Delta^{k}$). The ball $B_{\delta}(c_{v})\subset\mathbf{Y}$ of radius $\delta$ centered at $c_{v}$ covers $f^{-1}(S_{v})$. Consider an auxiliary map $\gamma:\Delta^{k}\to\mathbf{Y}$ defined by sending any vertex $v\in\Delta^{k}$ to $c_{v}\in\mathbf{Y}$, and then extending linearly on $\Delta^{k}$. Compose this map with $f$, and consider $\gamma\circ f:\mathcal{X}\to\mathbf{Y}$. Its image is a simplicial complex of dimension at most $k$, and to complete the proof it suffices to show that every point $x\in\mathcal{X}$ is moved by distance $\|x-\gamma(f(x))\|<\delta$. Let $f(x)$ lie in the relative interior of a simplex of $\Delta^{k}$ with the vertices $v_{0},\ldots,v_{m}$. Then $x\in f^{-1}(S_{v_{i}})\subset B_{\delta}(c_{v_{i}})$, and $\|x-c_{v_{i}}\|<\delta$, where $0\leq i\leq m$. By construction of $\gamma$, the point $\phi(f(x))$ lies in the convex hull of the points $c_{v_{i}}$, $0\leq i\leq m$. Therefore, $\|x-\gamma(f(x))\|$ does not exceed the maximum of $\|x-c_{v_{i}}\|$ over $0\leq i\leq m$, and this maximum is less than $\delta$. ∎
Remark A.2.
Depending on the geometry of $\mathbf{Y}$, this inequality may be slightly improved if for each vertex $f^{-1}(S_{v})$ we cover $f^{-1}(S_{v})$ by a ball of the smallest possible radius. For example, if $\mathbf{Y}=\mathbb{R}^{N}$ is Euclidean, then it follows from Jung’s theorem [Jun01] that this radius can be taken to be $\delta\sqrt{\frac{N}{2(N+1)}}$ (instead of $\delta$). The rest of the proof runs without changes, and the final result is $\operatorname{AW}_{k}(\mathcal{X}\subset\mathbb{R}^{N})\leq\sqrt{\frac{N}{2(N+1)}}\operatorname{UW}_{k}(\mathcal{X})<\frac{1}{\sqrt{2}}\operatorname{UW}_{k}(\mathcal{X})$. Another extreme example is when $\mathbf{Y}$ is hyperconvex (for example, $L^{\infty}(\mathcal{X})$), and any bounded set $\mathcal{A}$ can be covered by a ball of radius $\frac{1}{2}\operatorname{diam}(\mathcal{A})$. In this case, our estimates actually imply that $\operatorname{UW}_{k}(\mathcal{X})=2\operatorname{AW}_{k}(\mathcal{X}\subset\mathbf{Y})$.
The following property was used in Section 3.2.
Lemma A.3.
The treewidth enjoys the following continuity property. Let $\mathcal{X}_{1}\supset\mathcal{X}_{2}\supset\cdots$ be a nested sequence of compact sets in a Banach space $\mathbf{Y}$, and let $k$ be a non-negative integer. Then
$\lim_{i\to\infty}\operatorname{TW}_{k}(\mathcal{X}_{i}\subset\mathbf{Y})=\operatorname{TW}_{k}\left(\bigcap_{i}\mathcal{X}_{i}\subset\mathbf{Y}\right).$
Proof.
Denote $\mathcal{X}=\bigcap\mathcal{X}_{i}$, $w=\operatorname{TW}_{k}(\mathcal{X}\subset\mathbf{Y})$, and let $f:\mathcal{X}\to\mathbf{Y}$ be a witness map: a continuous map whose image lies in a finite contractible $k$-dimensional simplicial complex $\Delta$ and such that $\|x-f(x)\|<w+\epsilon/3$ for all $x\in\mathcal{X}$ and some arbitrarily chosen $\epsilon>0$. For each point $y\in\Delta$, the fiber $f^{-1}(y)$ lies in the open ball $B_{w+\epsilon/3}(y)$. Since $f$ is continuous on the compact set $\mathcal{X}$, it is uniformly continuous, hence there is a tiny radius $\rho>0$ such that such that the “thickened fiber” $f^{-1}(B_{\rho}(y))$ lies in the open ball $B_{w+2\epsilon/3}(y)$, for each $y\in\Delta$. Since $\Delta$ is a finite complex of dimension $k$, there is a finite closed cover $\bigcup D_{j}=\Delta$ such that its multiplicity is $k+1$, its nerve is homotopy equivalent to $\Delta$, each $D_{j}$ fits in a ball $B_{\rho}(y_{j})$, and $C_{j}:=f^{-1}(D_{j})$ fits in the ball $B_{w+2\epsilon/3}(y_{j})$.
We claim that the sets in the closed cover $\bigcup C_{j}=\mathcal{X}$ can be inflated a little bit while preserving their intersection pattern. For each collection of indices $J=\{j_{1},\ldots,j_{n}\}$ such that the intersection $\bigcap\limits_{j\in J}C_{j}$ is empty, it follows from the compactness of $\mathcal{X}$ that the number
$\delta_{J}=\min\limits_{x\in\mathcal{X}}\max_{j\in J}\operatorname{dist}(x,C_{j})$
is attained and positive. Take a positive $\delta$ smaller than $\epsilon/3$ and also smaller than each $\delta_{J}$ over all collections $J$ such that $\bigcap\limits_{j\in J}C_{j}=\varnothing$. Consider the open cover $\{U_{j}\}$ of $\mathcal{X}$, where $U_{j}=\mathcal{N}_{\epsilon}(C_{j}\subset\mathbf{Y})$. It has the same nerve as $\{C_{i}\}$, and each $U_{j}$ is contained in $B_{w+\epsilon}(y_{j})$. We have
$\mathcal{X}\subset\bigcup_{j}U_{j}\subset\mathcal{N}_{\epsilon}(\mathcal{X}\subset\mathbf{Y}).$
Now pick a partition of unity $\{\psi_{j}\}$ subordinate to $\{U_{j}\}$. Use it to map $\bigcup_{j}U_{j}$ to its nerve, which has the same dimension and homotopy type as $\Delta$. Namely, a point $x\in\bigcup_{j}U_{j}$ is mapped to $\sum\limits_{j}\psi_{j}(x)y_{j}$. Since $U_{j}\subset B_{w+\epsilon}(y_{j})$, it follows the distance from $x$ to its image is less than $w+\epsilon$.
Notice that the open set $\bigcup_{j}U_{j}$ contains all $\mathcal{X}_{i}$ for $i$ large enough, as it can be easily deduced from compactness. Therefore, we have a map from $\mathcal{X}_{i}$, for all $i$ large enough, to a contractible $k$-dimensional simplicial complex inside $\mathbf{Y}$, and every point is shifted by a distance less than $w+\epsilon$. Hence,
$\operatorname{TW}_{k}(\mathcal{X}_{i}\subset\mathbf{Y})-\operatorname{TW}_{k}(\mathcal{X}\subset\mathbf{Y})<\epsilon.$
Since $\epsilon$ was arbitrary, the result follows. ∎
Appendix B Übercontractible Sets
Here we only work in $\mathbf{Y}=\mathbb{R}^{N}$ with Euclidean metric, and we are interested in 00-übercontractible sets, which we call übercontractible. A rich source of übercontractible sets comes from the cut-locus construction.
Definition B.1.
Let $K\subset\mathbf{Y}$ be a convex body (that is, a compact convex set with a nonempty interior). For each $x\in K$, consider the largest closed ball centered at $x$ and contained in $K$; let $r^{\partial K}(x)\geq 0$ be its radius, and $B^{\partial K}(x)\subset\partial K$ be the set of the points where this ball touches the boundary of $K$. The cut-locus of $\partial K$ is the closure of the set of all such $x\in K$ for which the cardinality of $B^{\partial K}(x)$ is at least $2$:
$C^{\partial K}=\operatorname{cl}\big{(}\{x\in K:|B^{\partial K}(x)|>1\}\big{)}.$
There is a canonical retraction $c^{\partial K}:K\to C^{\partial K}$ defined as follows. For $x\in C^{\partial K}$, set $c^{\partial K}(x)=x$. For each point $x\in K\setminus C^{\partial K}$, let $b^{\partial K}(x)$ be the only element of $B^{\partial K}(x)$, and consider the largest closed ball contained in $K$ and touching $\partial K$ at $b^{\partial K}(x)$. The center of this ball belongs to the cut-locus and will be denoted $c^{\partial K}(x)$.
Theorem B.2.
The cut-locus $C^{\partial K}$ of the boundary of a convex body $K$ is übercontractible as long as it satisfies the following “tameness” assumption: $C^{\partial K}$ consists of finitely many compact convex sets.
The tameness assumption is introduced here in order to apply a lemma from Morse theory for the distance functional (B.3 below). It is highly likely that both the lemma and theorem can be stated and proven in greater generality, removing the tameness assumption completely. For now, we just remark that tame convex bodies in the sense of B.2 are $C^{2}$-dense among all convex bodies.
The key lemma in the proof is the following application of Morse theory for the distance functional. It was applied in [BHK15, Section 3] to give a sufficient condition for a set to be contractible, but the proof there, in fact, guarantees that the set is übercontractible.
Lemma B.3 (cf. [BHK15, Theorem 3.2]).
Let $\mathcal{T}\subset\mathbf{Y}$ be a union of finitely many compact convex sets. For each $x\notin\mathcal{T}$ consider the largest ball centered at $x$ whose interior does not meet $\mathcal{T}$. Let $B^{\mathcal{T}}(x)$ be the set of the points where this ball touches $\mathcal{T}$. Suppose that for each $x\notin\mathrm{conv}(B^{\mathcal{T}}(x))$. Then $\mathcal{T}$ is übercontractible.
Proof of B.2 for polytopes.
Suppose $K$ is a convex polytope. It can be shown that $C^{\partial K}$ is a polyhedral complex of codimension $1$ in $Y$. Let $F$ be a facet of $K$, and let $n_{F}$ be the corresponding inner normal vector. The (relative) boundary of $F$ lies in $C^{\partial K}$. The map $c^{\partial K}$, restricted to $F$, send each point $x\in F$ along $n_{F}$ until it hits $C^{\partial K}$. Together $F$ and the polyhedral surface $c^{\partial K}(F)$ bound a convex polytope $K_{F}$.
Consider a point $x\in K\setminus C^{\partial K}$. It then lies in $K_{F}$, for some facet $F$. It is sufficient to verify the assumption of B.3: one needs to check that $x\notin\mathrm{conv}(B^{\mathcal{T}}(x))$, where $\mathcal{T}=C^{\partial K}$. Indeed, for every $y\in B^{\mathcal{T}}(x)$, the vector $y-x$ forms an acute angle with $n_{F}$, and therefore, the entire set $B^{\mathcal{T}}(x)$ lies in the open halfspace $\{z\in Y:\langle z-x,n_{F}\rangle>0\}$. Therefore, $x\notin\mathrm{conv}(B^{\mathcal{T}}(x))$. ∎
Unfortunately, the case of the general convex body $K$ cannot be proven by approximating $K$ with polytopes $K_{i}$, because in order to guarantee $C^{0}$-convergence $C^{\partial K_{i}}\to C^{\partial K}$, we cannot get away with just polytopes. We sketch an argument that works for fairly general convex bodies, modulo some technicalities.
Sketch of the proof of B.2 for more general convex bodies, including tame ones.
Suppose our convex body $K$ can be approximated by a convex body $K^{\prime}$, satisfying the following three properties (these will be sufficient to prove the übercontractibility of $C^{\partial K}$).
  1. 1.
    The boundary of $K^{\prime}$ is $C^{2}$-smooth and strongly convex in the sense that its second fundamental form is positive definite.
  2. 2.
    The cut-locus of $\partial K^{\prime}$ consists of finitely many compact convex sets (we assume this to be able to use B.3).
  3. 3.
    The cut-locus of $\partial K^{\prime}$ approximates the cut-locus of $\partial K$ well enough so that the übercontractibility of the former implies the übercontractibility of the latter.
The third property is somewhat vague, so we start by arguing that the tameness assumption suffices. If $K$ is tame, we can inflate it slightly to get a new convex body $K^{\prime}$ that satisfies the first property, while preserving cut-locus $C^{\partial K^{\prime}}=C^{\partial K}=\mathcal{T}$ (which already satisfies the second property by the tameness assumption). This amounts to choosing $r^{\partial K^{\prime}}(x)$ for $x\in\mathcal{T}$ carefully, so that on every face of $\mathcal{T}$, $r^{\partial K^{\prime}}$ is $C^{1}$-close to $r^{\partial K}$, and $r^{\partial K^{\prime}}$ is strongly convex. The details are omitted.
In the rest of the proof we assume that $K$ satisfies the first two properties above, and we will verify the assumption of B.3 for $\mathcal{T}=C^{\partial K}$ to show that it is übercontractible. Pick points $x\in K\setminus\mathcal{T}$, and $y\in B^{\mathcal{T}}(x)$. Our goal is to prove that $\langle y-x,n\rangle>0$, where $n$ is the inner normal to $\partial K$ at $b^{\partial K}(x)$. This will imply that $x\notin\mathrm{conv}(B^{\mathcal{T}}(x))\subset\{z\in Y:\langle z-x,n\rangle>0\}$.
Suppose, for the contrary, that $\langle y-x,n\rangle\leq 0$. Parametrize the straight line segment $[x,y]$ linearly as $x(t),t\in[0,1]$, $x(0)=x$, $x(1)=y$. For each $t\in[0,1)$, let $u(t)=b^{\partial K}(x(t))$ be the only element $B^{\partial K}(x(t))$. Consider also $v(t)=c^{\partial K}(x(t))$, and notice that $v(1)=y$. The key idea is to look at the continuous family of straight line segments $[x(t),v(t)]$, and investigate how they intersect the open ball $O$ centered at $x$ of radius $r^{\partial K}(x)$. We make the following observations.
  • The open ball $O$ does not intersect $\mathcal{T}$, while $v(t)\in\mathcal{T}$. So for each $t\in[0,1)$, the segment $[x(t),v(t)]$ starts inside $O$ and ends outside of $O$.
  • The strong convexity of $\partial K$ implies that the angle between the vectors $y-x$ and $v(t)-x(t)$ is strictly increasing for $t\in[0,1)$. To see that, one needs to differentiate in $t$ the inner normal to $\partial K$ at $u(t)$ (which is collinear with $v(t)-x(t)$).
  • Initially, the angle between the vectors $y-x$ and $v(0)-x(0)$ is non-acute (since by assumption $\langle y-x,n\rangle\leq 0$). Therefore, for $t$ close to $1$ the angle between $y-x$ and $v(t)-x(t)$ is obtuse, and tends to a limit that is obtuse. It follows that the length of the part of $[x(t),v(t)]$ that lies in $O$ is bounded away from 00.
  • But, as $t\to 1$, the segment $[x(t),v(t)]$ degenerates to the point $y$, so its length must approach zero. This contradiction concludes the proof.
Appendix C Notation Table
Table 1: Notation and main symbols.
Notation Definition Place
$\mathrm{conv}(\mathcal{X})$ Convex hull of $\mathcal{X}\subset\mathbf{Y}$
$\overline{\mathrm{conv}}(\mathcal{X})$ Closure of convex hull of $\mathcal{X}\subset\mathbf{Y}$
$\mathrm{H}_{k}(\cdot)$ Singular homology in degree $k$
$\operatorname{\tilde{H}}_{k}(\cdot)$ Reduced singular homology in degree $k$
$\mathcal{N}_{r}(\mathcal{X}\subset\mathbf{Y})$ Open $r$-neighborhood of $\mathcal{X}$ in $\mathbf{Y}$ Section 2.1
$\overline{\mathcal{N}}_{r}(\mathcal{X}\subset\mathbf{Y})$ Closed $r$-neighborhood of $\mathcal{X}$ in $\mathbf{Y}$ Section 2.1
$\mathcal{V}_{r}(\mathcal{X})$ Vietoris-Rips complex of $\mathcal{X}$ for distance $r$ 2.1
$\mathcal{C}_{r}(\mathcal{X}\subset\mathbf{Y})$ech complex of $\mathcal{X}\subset\mathbf{Y}$ for distance $r$ 2.2
$\mathrm{Spec}_{k}(\Delta_{\bullet})$ Homological spectrum of the filtration $\Delta_{\bullet}$ 2.17
$\mathrm{PD}_{k}(\mathcal{X})$ $k^{th}$-persistence diagram of $\mathcal{X}$ with VR filtration Section 2.1.2
$\widecheck{\mathrm{PD}}_{k}(\mathcal{X}\subset\mathbf{Y})$ $k^{th}$-persistence diagram of $\mathcal{X}\subset\mathbf{Y}$ with Čech filtration Section 2.1.2
$b_{\omega},d_{\omega}$ Birth and death time of the homology class $\omega$ 2.18
$d_{\omega}-b_{\omega}$ Persistence (lifespan) of the homology class $\omega$ 2.18
$\mathbf{d}_{b}(\cdot,\cdot)$ Bottleneck distance between persistence diagrams Section 2.1.3
$\mathbf{d}_{\mathrm{H}}^{\mathbf{Y}}(\cdot,\cdot)$ Hausdorff distance between two subsets in $\mathbf{Y}$ Section 2.1
$\mathbf{d}_{\mathrm{GH}}(\cdot,\cdot)$ Gromov–Hausdorff distance between two metric spaces Section 2.1
$L^{\infty}(\mathcal{X})$ Space of bounded functions on $\mathcal{X}$ with $\sup$ norm Section 2.2
$\rho(\mathcal{M})$ Gromov’s filling radius of $\mathcal{M}$ 2.30
$\rho(\omega;\mathcal{X}\subset\mathbf{Y})$ The relative filling radius of $\omega\in\operatorname{\tilde{H}}_{k}(\mathcal{X})$ relative to $\mathbf{Y}$ 2.31
$\rho(\omega;\mathcal{X})$ The absolute filling radius of $\omega\in\operatorname{\tilde{H}}_{k}(\mathcal{X})$ 2.31
$\mathrm{rad}(\mathcal{X})$ The radius of $\mathcal{X}$ 2.35
$\mathrm{rad}(\mathcal{X}\subset\mathbf{Y})$ The circumradius of $\mathcal{X}$ relative to $\mathbf{Y}$ 2.35
$\operatorname{UW}_{k}(\mathcal{X})$ Urysohn $k$-width of $\mathcal{X}$ 3.1
$\operatorname{AW}_{k}(\mathcal{X}\subset\mathbf{Y})$ Alexandrov $k$-width of $\mathcal{X}$ relative to $\mathbf{Y}$ 3.4
[…]
$\operatorname{TW}^{C}_{k}(\mathcal{X}\subset\mathbf{Y})$ $C$-robust $k$-dimensional treewidth of $\mathcal{X}$ 3.19
$\operatorname{KW}_{k}(\mathcal{X}\subset\mathbf{Y})$ Kolmogorov $k$-width of $\mathcal{X}$ relative to $\mathbf{Y}$ 3.6
[…]
$\operatorname{{\ddot{u}}-spread}(\mathcal{X}\subset\mathbf{Y})$berspread of $\mathcal{X}$ relative to $\mathbf{Y}$ 4.4
$\xi(\mathcal{X})$ VR-extinction time of $\mathcal{X}$ 5.1
$\widecheck{\xi}(\mathcal{X}\subset\mathbf{Y})$ech-extinction time of $\mathcal{X}$ relative to $\mathbf{Y}$ 5.1
$\operatorname{cdef}(\mathcal{X}\subset\mathbf{Y})$ Convexity deficiency of $\mathcal{X}$ relative to $\mathbf{Y}$ 5.4
$\operatorname{E}(\mathcal{X})$ Tight span of $\mathcal{X}$ 2.27
$\operatorname{hcdef}(\mathcal{X})$ Hyperconvexity deficiency of $\mathcal{X}$ 5.8